### v4

# Bayes, E-values, and Testing

Abstract
We separate sequential evidence into three typed layers: (i) representation (Radon–Nikodým / likelihood-ratio geometry), (ii) validity (Ville/martingale control under optional stopping), and (iii) decision (boundary choice and power/efficiency calibration). Within the coherent predictive/log-loss subclass, the Bayes-risk/Fubini decomposition forces a likelihood-ratio (Bayes-factor) representation of evidence, while validity-only e-process constructions remain strictly broader. A typed calculus then organizes these layers and their bridges: the monoidal log-loss map connects Bayes factors, sequential testing, and information-theoretic regret through Good’s weight of evidence; Markov/Ville inequalities supply anytime-valid certificates; and decision-theoretic cutoffs determine rejection regions from priors and losses rather than from Markov bounds. Within the coherent predictive subclass, this reduces evidence to likelihood-ratio or Bayes-factor testing; general e-process constructions may extend beyond this representation. We separate representation from construction: expectation-based recipes (e.g. Markov cutoffs) produce valid certificates but need not recover Bayes-optimal rejection regions, whereas the Fubini decomposition yields likelihood-ratio regions directly from Bayes risk. A typing discipline formalizes when bridges exist between probabilistic evidence, coding/MDL objects, and algorithmic randomness—and when they provably fail: NML codes violate filtration-measurability, while prequential codes [11] and the universal semimeasure provide valid but non-computable alternatives.
Keywords: E-values; Bayes factors; supermartingales; KL divergence; Sanov theorem; optional stopping; computational boundary; conformal prediction; weight of evidence; game-theoretic probability.
1 Introduction
1.1 The evidence problem
Statistical evidence takes multiple formally distinct forms: $p$-values, likelihood ratios, Bayes factors, E-values, predictive log-loss, and description length. A $p$-value is a tail probability under a fixed null; a Bayes factor is a marginal likelihood ratio; an E-value is a nonnegative random variable with $\operatorname{\mathbb{E}}_{H_{0}}[E]\leq 1$; a code length requires computability assumptions. Conflating these objects can lead to misapplied composition rules, stopping guarantees unsupported without additional conditions, and tempting but non-equivalent identifications: $p$-values used sequentially without correction (conflating laws with processes), code lengths equated with log-probabilities without semimeasure conditions (conflating codes with laws), and conformal certificates confused with predictive distributions.
Central thesis.
This paper develops a framework for sequential evidence in which probability, prediction, sequential testing, and coding are separated into four formally distinct levels (Section 1.2). The negative log-probability $\ell_{P}(x^{n}):=-\log P(x^{n})$ acts as a monoidal bridge from multiplicative probability to additive evidence; KL divergence governs both evidence growth rates and large-deviation rarity; and a computational boundary theorem delineates where code lengths yield valid E-processes and where they provably fail. A log-loss/Fubini decomposition of Bayes risk (Section 2.3) shows that the likelihood ratio is the canonical bridge from predictive laws to sequential evidence when evidence is required to arise from coherent prediction; Markov’s inequality then supplies universal validity certificates, while likelihood-ratio structure provides the Bayes-risk-aligned semantics. Thus, within the coherent predictive subclass, the likelihood ratio is the canonical evidence process under log-loss. In particular, for composite hypotheses the Bayes predictive likelihood ratio $\Lambda_{n}$ provides a canonical E-variable $E_{n}=\Lambda_{n}$ within the mixture/predictive regime, highlighting likelihood ratios as a distinguished representation of sequential evidence when tied to coherent prediction under log-loss. A central message is that inequality-calibrated evidence thresholds (e.g. Markov/Ville cutoffs) guarantee validity but can conflict with Bayes-risk optimality: validity alone does not select the critical region that optimizes Type I/Type II tradeoffs.
Nothing here disputes the Ville-validity of e-processes or their utility as stopping-safe certificates. Our focus is the representational interface: when evidence is required to arise from coherent prediction under log-loss, LR/BF structure is forced, and boundary design becomes an efficiency problem rather than a validity problem.
Likelihood ratio as the canonical evidence primitive.
[…]
Validity is not optimality.
E-value methods often begin from the constraint $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ and appeal to Markov’s inequality to obtain the anytime-valid rule $\operatorname{\mathbb{P}}_{0}(E\geq c)\leq 1/c$. This delivers universal validity, but it does not identify the Bayes-risk-optimal rejection region. When the validity certificate (Markov/Ville) is not distinguished from the representation of evidence, the likelihood ratio can appear dispensable. Our analysis clarifies the distinction: Markov/Ville calibration certifies error control; the likelihood ratio supplies the Bayes-risk-aligned critical region and semantics once evidence is tied to prediction under log-loss. Our contribution separates two layers that are sometimes not distinguished: (i) the validity certificate provided by martingale inequalities, and (ii) the efficiency of boundary selection when additional likelihood-ratio structure is present. Boundary choice lives in the execution/decision layer and should be tuned to the intended risk-efficiency criterion once a representational model is fixed.
Representation versus construction within the typed calculus.
[…]
1.2 A typed solution: four levels of discourse
[…]
We distinguish four levels of mathematical objects. (L1) Sequences: finite strings $x^{n}=(x_{1},\dots,x_{n})\in\mathcal{X}^{n}$ and infinite sequences $x^{\infty}\in\mathcal{X}^{\infty}$. (L2) Laws / predictive specifications: probability measures $P$ on $(\mathcal{X}^{\infty},\mathcal{F})$ or predictive kernels $p(\cdot\,|\,x^{t-1})$, linked by the chain rule $P(x^{n})=\prod_{t=1}^{n}p(x_{t}\,|\,x^{t-1})$. (L3) Processes on filtrations: filtered spaces $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\operatorname{\mathbb{P}})$ and nonnegative supermartingales adapted to $(\mathcal{F}_{t})$; optional stopping, anytime validity, and sequential error control live at this level. (L4) Codes / descriptions: prefix-free codes, Kolmogorov complexity $K(x)$, MDL/NML code lengths; the bridge to (L2) is $K(x)=-\log m(x)+O(1)$ for the universal semimeasure $m$. Throughout, we explicitly distinguish objects by their level and only apply bridges when formally justified. These distinctions prevent common confusions, for example between stopping-rule dependent $p$-values and stopping-rule invariant likelihood-ratio evidence. Each theorem below declares its level (L1–L4) and the bridge(s) it invokes; equalities such as $\ell_{P}(x^{n})=-\log P(x^{n})$ are bridges between levels, not identifications.
1.3 Main contributions
The paper establishes six structural results.
(C1) Bayes-risk decomposition and LR representation. We derive a Fubini/tower decomposition of Bayes risk under log-loss that yields Good’s weight of evidence and identifies the (mixture) likelihood ratio $B_{t}=dM_{1}/dM_{0}|\mathcal{F}_{t}$ as the canonical E-process within the coherent predictive subclass (Theorem˜2.6); the corresponding critical region is decision-theoretic (priors and losses), not inequality-chosen (Section˜3.7). In this sense, within the log-loss/Bayes-risk coherent subclass, sequential evidence admits a likelihood-ratio representation; outside this subclass, e-processes may exist without such a representation. For uniform validity over arbitrary composite null classes $\mathcal{P}$ (i.e., $\operatorname{\mathbb{E}}_{P}[E_{t}]\leq 1$ for every $P\in\mathcal{P}$), Larsson et al. [27] develop the numeraire e-variable via reverse information projection, which always exists without any assumptions on $\mathcal{P}$.
(C2) Monoidal bridge. The log-score map $\mathcal{L}:P\mapsto\ell_{P}=-\log P$ translates multiplicative probability into additive evidence, sending products to sums and mixtures to log-sum-exp (Proposition˜2.3). This bridge is the structural foundation linking Bayesian, information-theoretic, and coding-theoretic formulations of evidence.
(C3) E-process algebra. Bayes factors and likelihood-ratio processes are nonnegative supermartingales under the null, and the evidence class $\mathcal{E}(H_{0})$ is closed under convex mixtures, products under conditional independence, and optional stopping, with anytime validity guaranteed by Ville’s inequality (Theorems˜4.5 and 4.3). The sequential LR martingale admits a symmetry: $B_{t}$ is a $P_{0}$-martingale while $1/B_{t}$ is a $P_{1}$-martingale.
(C4) KL rate theory. The growth rate $(1/t)\log E_{t}\to D_{\operatorname{KL}}(P_{1}\|P_{0})$ a.s. under correct specification, and becomes $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})$ under misspecification (Propositions˜5.4, 5.3 and 5.7). Sanov and inverse Sanov are dual manifestations of the same rate function with reversed KL arguments.
(C5) Computational boundary. Code lengths do not automatically yield E-processes (Proposition˜7.3). The passage from (L4) to (L3) requires a genuine probability measure or the universal semimeasure; NML/MDL codes fail because their normalizing constants depend on the full sample size, violating the filtration-measurability required for supermartingale structure. Prequential codes succeed (Section˜7.5).
(C6) Exchangeability testing. Conformal e-prediction (Proposition˜6.3) constructs E-values valid under permutation invariance, closing the loop between symmetry, evidence, and sequential validity.
Reader orientation.
Under log-loss, the Fubini/Bayes-risk decomposition singles out the likelihood ratio as the canonical evidence representation (C1). Markov/Ville inequalities provide universal validity certificates, but they operate at a polynomial $1/b$ scale and do not determine the optimal critical region. Confusing validity with optimality can lead to boundaries that sacrifice efficiency: the Bayes-risk-optimal boundary is governed by the moderate-deviation scale $\sqrt{(\log b)/t}$ of the log-likelihood ratio, not by Markov’s inequality (Remarks˜2.7 and 3.8). The paper’s central contribution is to separate representation (LR from Bayes risk) from construction (any valid E-process), and to show that the asymptotic geometry of these two layers differs fundamentally.
Positioning.
We build on the E-value framework [42, 21], game-theoretic probability [36], Dawid’s prequential principle [11, 12], Goldstein’s prevision framework [17], and the MDL programme [35, 20]. We identify the precise structural conditions under which these frameworks coincide and where they provably diverge. Our results complement the safe-testing perspective by isolating structural conditions under which E-process constructions succeed or fail [22].
Scope relative to the e-process literature.
In modern terminology, e-processes are defined by optional-stopping control and strictly generalize test (super)martingales; in particular, product constructions from sequential e-values yield valid e-processes but do not exhaust the class [33, Def. 7.3]. Our goal is different: we identify the typed interfaces under which evidential objects admit a coherent predictive interpretation, and show that within this coherent subclass the likelihood ratio (or a mixture of likelihood ratios) is the canonical representation of evidence, while many construction-level recipes remain valid yet non-optimal from a Bayes-risk standpoint. For composite alternatives, selecting a mixture model turns the Bayes factor into a likelihood ratio against that mixture; hence the Bayes factor is itself an E-value under the mixture (prior-predictive) null (Proposition˜3.9), and the remaining subtlety is not validity but execution: the cutoff must be specified decision-theoretically (Section˜3.7) rather than inherited from Markov-style bounds. General e-process constructions and their distinction from test supermartingales are treated in detail in the recent e-process literature (e.g., Ramdas 33, Ch. 7); our focus is the Bayes-risk/log-loss route that singles out LR/BF as canonical within the coherent predictive subclass.
1.4 Structural roadmap
The paper follows a logical arc from objects to morphisms to invariants to boundaries:
  1. 1.
    Probability $\to$ log-loss bridge (Section 2). The monoidal bridge $\ell_{P}=-\log P$ and Good’s weight of evidence.
  2. 2.
    Log-loss differences $\to$ E-processes (Section 3). Likelihood-ratio and mixture constructions on filtered spaces; Ville’s inequality as the compositional guarantee.
  3. 3.
    Composition algebra (Section 4). The evidence class, convex closure, stitching, Gibbs amalgamation, and typed obstructions.
  4. 4.
    KL divergence $\to$ large deviations and evidence growth (Section 5). Sanov/inverse-Sanov duality; Gibbs variational characterization; information-regularized inference.
  5. 5.
    Symmetry $\to$ exchangeability testing (Section 6). De Finetti’s theorem and conformal e-prediction as invariance testing.
  6. 6.
    Coding $\to$ computational boundary (Section 7). Kolmogorov complexity, MDL, prequential codes, and the formal boundary separating computational from probabilistic claims.
Each section identifies the typed level of its objects, the bridges it invokes, and the non-bridges that constrain its scope. Figure˜1 summarizes the typed bridges and their admissible compositions.
1.5 Main results
We state informally the three flagship results of the paper. Full statements appear in the sections indicated.
Theorem 1.1 (Evidence class algebra; informal).
Let $H_{0}$ be a null hypothesis on a filtered space $(\Omega,\mathcal{F},(\mathcal{F}_{t}),\operatorname{\mathbb{P}})$. Define
$\mathcal{E}^{\mathrm{proc}}(H_{0}):=\bigl\{(E_{t})_{t\geq 0}:E_{t}\geq 0,\;(E_{t})\text{ supermartingale under }H_{0},\;\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1\bigr\}.$
Then $\mathcal{E}^{\mathrm{proc}}(H_{0})$ satisfies:
  1. (a)
    Convex mixtures (Proposition˜4.3): $w_{i}\geq 0$, $\sum w_{i}=1$ implies $\bigl(\sum w_{i}E^{i}_{t}\bigr)\in\mathcal{E}^{\mathrm{proc}}(H_{0})$.
  2. (b)
    Bayesian mixtures (Proposition˜4.9): $\int E^{\theta}_{t}\,\pi(d\theta)\in\mathcal{E}^{\mathrm{proc}}(H_{0})$.
  3. (c)
    Stopping (Theorem˜4.5): $E_{\tau}\in\mathcal{E}(H_{0})$ and $\operatorname{\mathbb{P}}_{H_{0}}(\sup_{t}E_{t}\geq c)\leq 1/c$.
Proposition 1.2 (KL evidence growth rate; informal).
Let $E_{n}=\prod_{t=1}^{n}P_{1}(X_{t})/P_{0}(X_{t})$ be the likelihood-ratio E-process.
  1. (a)
    Under correct specification ($X_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P_{1}$):
    $\frac{1}{n}\log E_{n}\xrightarrow{\mathrm{a.s.}}D_{\operatorname{KL}}(P_{1}\|P_{0}).$
  2. (b)
    Under misspecification ($X_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P_{\mathrm{true}}\neq P_{1}$):
    $\frac{1}{n}\log E_{n}\xrightarrow{\mathrm{a.s.}}D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1}).$
    The E-process drifts downward when $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})>D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})$.
Full statement: Proposition˜5.4. Connection to Sanov: Theorems˜5.3 and 5.7.
Proposition 1.3 (Computational boundary; informal).
Let $\mathcal{C}$ be a coding scheme and define $E_{n}^{\mathcal{C}}:=P_{0}(X^{n})^{-1}\cdot 2^{-\ell_{\mathcal{C}}(X^{n})}$.
  1. (a)
    If $\mathcal{C}$ corresponds to a probability measure $Q$, then $E_{n}^{\mathcal{C}}$ is a valid E-process.
  2. (b)
    If $\mathcal{C}$ is the NML code, then $E_{n}^{\mathcal{C}}$ is not in general a supermartingale: NML’s normalizing constant depends on the full sample size $n$, so sequential factors are not valid predictive kernels.
  3. (c)
    If $\mathcal{C}$ induces the universal semimeasure $m$, then $m(X^{n})/P_{0}(X^{n})$ is a valid but non-computable E-process.
Full statement: Proposition˜7.3. Prequential repair: Section˜7.5.
1.6 Notation and conventions
Logs and units.
Unless stated otherwise, $\log$ denotes the natural logarithm (base $e$); quantities are therefore measured in nats. When interpreting code lengths or evidence in bits, we write $\log_{2}$ explicitly (equivalently, multiply nats by $1/\ln 2$). KL divergences, entropy rates, and evidence growth rates all use natural logarithms unless otherwise noted.
Filtration.
$\mathcal{F}_{t}:=\sigma(X_{1},\dots,X_{t})$ unless otherwise specified.
Divergence.
$D_{\operatorname{KL}}(P\|Q):=\int\log(dP/dQ)\,dP$ when $P\ll Q$, and $\infty$ otherwise.
1.7 A worked example: comparing evidence measures
Consider a concrete problem: we observe $n=20$ coin flips with $k=14$ heads and wish to evaluate evidence regarding the null hypothesis $H_{0}:p=0.5$ (fair coin) against the alternative $H_{1}:p>0.5$.
The p-value (L2/L3: tail probability; depends on the stopping rule / look policy) is the probability under $H_{0}$ of observing a result at least as extreme:
$p\text{-value}=\operatorname{\mathbb{P}}(X\geq 14\mid H_{0})=\sum_{i=14}^{20}\binom{20}{i}(1/2)^{20}\approx 0.058.$
At the conventional $\alpha=0.05$ threshold, this narrowly misses significance.
The Bayes factor (L2: marginal likelihood ratio) compares marginal likelihoods. With a uniform prior on $p\in(0.5,1]$ under $H_{1}$:
$\text{BF}_{10}=\frac{2\binom{20}{14}\int_{0.5}^{1}p^{14}(1-p)^{6}\,dp}{\binom{20}{14}(1/2)^{20}}\approx 2.38.$
On Jeffreys’ scale, this falls in the “anecdotal” range.
The likelihood-ratio E-value at the MLE (L3: wealth factor; supermartingale payoff under the null if used sequentially) with $\hat{p}=0.7$:
$E=\left(\frac{0.7}{0.5}\right)^{14}\left(\frac{0.3}{0.5}\right)^{6}\approx 5.18.$
A sequential bettor wagering against the fair-coin hypothesis would have multiplied their stake by a factor of 5.18, with anytime validity guaranteed by Ville’s inequality regardless of the stopping rule.
All three measures agree qualitatively: the data lean toward the coin being biased, but the evidence is not yet strong enough to be conclusive. The differences in magnitude reflect their different typed guarantees, frequentist tail probability (L2/L3), Bayesian likelihood ratio (L2), and game-theoretic wealth (L3), rather than disagreement about the data.
The Edwards–Lindman–Savage die. A classic illustration, due to Edwards et al. [15], sharpens the contrast between significance and evidence. Suppose we test whether a particular face of a die has probability $p_{0}=1/6$ (fair) versus $p_{1}=1/5$ (slightly loaded) based on $n$ independent rolls. The likelihood ratio after observing $k$ target faces in $n$ rolls is
$\mathrm{LR}_{n}\;=\;\Bigl(\frac{6}{5}\Bigr)^{\!k}\Bigl(\frac{24}{25}\Bigr)^{\!n-k}.$
Because the two hypotheses are close—the Kullback–Leibler divergence $D_{\operatorname{KL}}(p_{1}\,\|\,p_{0})$ is only about $0.004$ nats—evidence accumulates slowly: even after n=1,000n=1{,}000 rolls the expected log-likelihood ratio under $H_{1}$ is roughly $3.8$ nats ($\approx 5.5$ bits), a Bayes factor of about $45$. Meanwhile a two-sided $z$-test rejects $H_{0}$ at level $0.05$ well before such evidence accrues, illustrating that frequentist significance can diverge sharply from Bayesian weight of evidence when the alternative is close to the null. Lopes and Polson [30] revisit this example in a Bayesian hypothesis testing framework; see also Ram [1, Example 8.1] for a game-theoretic reinterpretation as an e-value.
2 The evidence functional as a monoidal bridge
The map $P\mapsto-\log P$ is a monoidal bridge: it translates multiplicative probability into additive evidence, code lengths, and cumulative loss.
2.1 Additive evidence from probability
Definition 2.1 (Log-score map).
The log-score map is the functor $\mathcal{L}:P\mapsto-\log P$ that sends a probability measure (or predictive kernel) to its pointwise negative logarithm. Applied to a data realization $x^{n}$, $\mathcal{L}(P)(x^{n})=-\log P(x^{n})$. This map converts multiplicative structure (products of probabilities) to additive structure (sums of log-scores), and is the canonical bridge from level (L2) to the evidence and coding scales at levels (L1) and (L4). Theorem˜B.2 shows that this choice is structurally forced: the negative logarithm is the unique continuous order-preserving transform satisfying the multiplicative-to-additive requirement.
Definition 2.2 (Loss / surprisal / code length functional).
Given a law or predictive specification $P$ (an object at level L2), define the additive evidence (loss) for a data realization $x^{n}$ (an object at level L1) as
$\ell_{P}(x^{n}):=-\log P(x^{n}).$
When $P$ is specified via a predictive kernel $p(\cdot\,|\,x^{t-1})$, the chain rule gives
$\ell_{P}(x^{n})=\sum_{t=1}^{n}\bigl[-\log p(x_{t}\,|\,x^{t-1})\bigr],$
expressing the total loss as the sum of sequential surprisals.
This functional appears in multiple guises. In Shannon’s noiseless coding theorem, $\ell_{P}(x^{n})$ is the ideal code length for $x^{n}$ under the distribution $P$ [37]. In sequential prediction, $\ell_{P}(x^{n})$ is the cumulative log-loss incurred by a forecaster who assigns predictive probabilities according to $P$ [9]. In Bayesian inference, $\ell_{P}(x^{n})=-\log P(x^{n}\mid M)$ is the negative log-marginal likelihood, measuring how “surprising” the data are under model $M$.
Proposition 2.3 (Log is a monoidal bridge: product $\mapsto$ sum, mixture $\mapsto$ log-sum-exp).
[…]
Let $P,Q$ be probabilistic objects on compatible measurable spaces.
  1. 1.
    (Product $\mapsto$ sum.) For independent concatenation under the product measure $P\otimes Q$,
    $\ell_{P\otimes Q}(x^{n},y^{m})=\ell_{P}(x^{n})+\ell_{Q}(y^{m}).$
  2. 2.
    (Mixture $\mapsto$ log-sum-exp.) For a mixture $\bar{P}:=\int P_{\theta}\,\pi(d\theta)$,
    $\ell_{\bar{P}}(x^{n})=-\log\int\exp\bigl(-\ell_{P_{\theta}}(x^{n})\bigr)\,\pi(d\theta).$
Proof.
Part (1) follows from $P\otimes Q(x^{n},y^{m})=P(x^{n})\cdot Q(y^{m})$ and the logarithm converting products to sums. Part (2) is the definition of mixture applied under the logarithm: $\bar{P}(x^{n})=\int P_{\theta}(x^{n})\,\pi(d\theta)$, so $\log\bar{P}(x^{n})=-\log\int\exp(-\ell_{P_{\theta}}(x^{n}))\,\pi(d\theta)$. ∎
Remark 2.4.
Property (1) underlies the combination of independent E-values by multiplication; property (2) underlies Bayes factors as integrated likelihood ratios.
2.2 Good’s weight of evidence
Definition 2.5 (Weight of evidence).
For two models or hypotheses $H_{1},H_{0}$ with marginal or predictive probabilities $P_{1},P_{0}$, Good’s weight of evidence is
$W(x^{n}):=\log\frac{P_{1}(x^{n})}{P_{0}(x^{n})}=\ell_{P_{0}}(x^{n})-\ell_{P_{1}}(x^{n}).$
Positive weight indicates evidence for $H_{1}$; negative for $H_{0}$. Good [18] showed this is the unique measure both additive across independent observations (Proposition˜2.3) and consistent with the likelihood principle. Its exponential $\exp(W(x^{n}))=P_{1}(x^{n})/P_{0}(x^{n})$ is the canonical E-value.
The connection to Bayesian updating is immediate. If the prior odds are $\operatorname{\mathbb{P}}(H_{1})/\operatorname{\mathbb{P}}(H_{0})$, then the posterior odds satisfy
$\log\frac{\operatorname{\mathbb{P}}(H_{1}\,|\,x^{n})}{\operatorname{\mathbb{P}}(H_{0}\,|\,x^{n})}=\log\frac{\operatorname{\mathbb{P}}(H_{1})}{\operatorname{\mathbb{P}}(H_{0})}+W(x^{n}),$
so the weight of evidence is precisely the additive update from prior to posterior on the log-odds scale.
2.3 Bayes risk, Fubini, and weight of evidence
Under log-loss, the Bayes risk of a predictive specification $P$ is the expected cumulative loss $\operatorname{\mathbb{E}}[\ell_{P}(X^{n})]$, where $\ell_{P}(x^{n}):=-\log P(x^{n})$ (Section 2). For two specifications $P_{0}$ (null) and $P_{1}$ (alternative), the log-loss difference is
$\ell_{P_{0}}(x^{n})-\ell_{P_{1}}(x^{n})\;=\;\log\frac{P_{1}(x^{n})}{P_{0}(x^{n})},$
i.e. Good’s weight of evidence (Definition˜2.5). When $P_{i}$ are given via one-step predictive kernels, the tower property yields the sequential decomposition
$\log\frac{P_{1}(X^{n})}{P_{0}(X^{n})}\;=\;\sum_{t=1}^{n}\log\frac{p_{1}(X_{t}\mid X^{t-1})}{p_{0}(X_{t}\mid X^{t-1})}.$
[…]
Theorem 2.6 (Canonical evidence under log-loss).
Consider testing $H_{0}$ versus $H_{1}$, where either hypothesis may be composite. Let $\Pi_{0},\Pi_{1}$ be priors over the respective model classes and define the prior-predictive mixtures
$M_{0}^{(n)}(\cdot)=\int P_{F}^{\otimes n}(\cdot)\,\Pi_{0}(dF),\quad M_{1}^{(n)}(\cdot)=\int P_{F}^{\otimes n}(\cdot)\,\Pi_{1}(dF).$
Assume mutual absolute continuity. Under log-loss Bayes risk, the Bayes-risk-optimal test is a threshold rule in
$\Lambda_{n}(X^{n}):=\frac{dM_{1}^{(n)}}{dM_{0}^{(n)}}(X^{n}),$
and the canonical multiplicative evidence is $E_{n}=\Lambda_{n}$, with additive evidence $W_{n}=\log\Lambda_{n}$.
Sketch.
Under log-loss, Bayes risk can be written as an expectation of cumulative log-loss differences. Applying Fubini/Tonelli swaps the prior and data integrals and yields a pointwise posterior decision rule. The resulting optimal rejection region is of the form $\Lambda_{n}(X^{n})>\tau$ for a threshold $\tau$ depending on prior weights and losses. ∎
2.3.1 Fubini, Bayes risk, and the critical region
Consider binary action $\delta_{t}\in\{0,1\}$ at (optional) time $t$ with losses $L_{10}$ (reject $H_{0}$ when true) and $L_{01}$ (fail to reject when $H_{1}$ true), and prior weights $\pi_{0},\pi_{1}$. Fubini reduces Bayes risk to a pointwise posterior decision: reject $H_{0}$ whenever posterior odds exceed the loss ratio, i.e.
$\frac{\pi_{1}}{\pi_{0}}\,B_{t}\;\geq\;\frac{L_{10}}{L_{01}},\qquad\text{equivalently}\qquad B_{t}\geq c^{\star}:=\frac{\pi_{0}L_{10}}{\pi_{1}L_{01}}.$
Thus the critical region is determined by posterior odds and losses. Inequality-based bounds (e.g. Markov/Ville) certify error control for a chosen $c^{\star}$, but do not determine the optimal cutoff itself. A closely related Fubini/Bayes-risk route to LR optimality appears in Datta et al. [10]. The full decision-theoretic development is in Section˜3.7.
2.3.2 Two Bayes-risk decompositions and the role of the cutoff
Bayes risk admits two complementary decompositions.
(i) Posterior / critical-region form.
Let $\delta_{t}(x^{t})\in\{0,1\}$ be a decision rule at time $t$ and write $B_{t}(x^{t}):=\frac{\pi_{1}}{\pi_{0}}\frac{dM_{1}}{dM_{0}}\big|_{\mathcal{F}_{t}}(x^{t})$ for the posterior odds. By the tower property and Fubini,
$\mathrm{BR}(\delta_{t})=\operatorname{\mathbb{E}}\!\Big[\operatorname{\mathbb{E}}\!\big[L(\delta_{t},\theta)\mid\mathcal{F}_{t}\big]\Big]=\operatorname{\mathbb{E}}\!\Big[\min_{\delta\in\{0,1\}}\operatorname{\mathbb{E}}\!\big[L(\delta,\theta)\mid\mathcal{F}_{t}\big]\Big],$
so pointwise optimality holds inside the outer expectation, yielding a posterior-threshold rule: the optimal decision is a critical region $\{B_{t}\geq c_{t}\}$, where $c_{t}$ is determined by priors and losses. For composite hypotheses, the admissible Bayes-optimal E-process is the mixture likelihood ratio $B_{t}=dM_{1}/dM_{0}|_{\mathcal{F}_{t}}$, i.e. $E_{t}=\Lambda_{t}$ in the evidence layer (Theorem˜2.6).
(ii) Frequentist Type I/Type II form.
Equivalently, Bayes risk can be written as a prior-weighted average of Type I and Type II errors:
$\mathrm{BR}(\delta_{t})=\pi_{0}L_{10}\,\alpha_{t}(\delta_{t})+\pi_{1}L_{01}\,\beta_{t}(\delta_{t}),$
where $\alpha_{t}=\operatorname{\mathbb{P}}_{0}(\text{reject at }t)$ and $\beta_{t}=\operatorname{\mathbb{P}}_{1}(\text{accept at }t)$. This form makes explicit that choosing $c_{t}$ is choosing a point on the Type I/Type II frontier.
Choosing $c_{t}$ (execution-theoretic).
For fixed $(\pi_{0},\pi_{1},L_{10},L_{01})$, the Bayes rule yields a constant boundary $c^{\star}=\pi_{0}L_{10}/(\pi_{1}L_{01})$ in posterior odds (hence in $B_{t}$). In sequential monitoring, one may also impose operational constraints (e.g. anytime Type I control), which introduces a family of feasible boundaries $c_{t}$. Crucially, Markov/Ville inequalities only certify validity for a chosen boundary; they do not optimize it. Optimization of $c_{t}$ is governed by likelihood-ratio growth and moderate-deviation behavior of $\log B_{t}$, which controls the Type I/Type II tradeoff along the frontier. Accordingly, procedures based solely on Markov/Ville calibration are typically not Pareto-optimal in the $(\alpha_{t},\beta_{t})$ plane when likelihood-ratio structure is available: one can often reduce Type II error at the same Type I level (or vice versa) by using LR/Bayes-risk boundaries. In short: Markov/Ville choose a boundary after the fact (exogenously, for validity); Bayes risk chooses it endogenously (from priors, losses, and the LR geometry).
Remark 2.7 (Asymptotic boundary: $1/b$ versus $\sqrt{(\log b)/t}$).
Two asymptotic regimes govern sequential evidence thresholds, and confusing them obscures the role of the critical region.
  1. (i)
    Markov/Ville (validity). The universal bound $\operatorname{\mathbb{P}}_{0}(\sup_{s\leq t}E_{s}\geq b)\leq 1/b$ operates on $E_{t}$ at a polynomial $\sim\!1/b$ scale. It guarantees Type I control but says nothing about power or Bayes-risk optimality.
  2. (ii)
    Likelihood-ratio / moderate deviations (optimality). Under the alternative, $(1/t)\log B_{t}\to D_{\operatorname{KL}}(P_{1}\|P_{0})$ a.s. Moderate-deviation theory shows that the fluctuations of $(1/t)\log B_{t}$ around its KL slope are of order $\sqrt{(\log b)/t}$, an exponentially sharper scale that governs boundary selection and the Type I/Type II frontier.
Here $b$ denotes a chosen evidence threshold (e.g. $b=1/\alpha$); the appearance of $\log b$ reflects the conversion of a boundary crossing $B_{t}\geq b$ to the log-scale condition $\log B_{t}\geq\log b$, not an independent asymptotic variable. Intrinsic time-asymptotics are expressed in terms of $t$ (or $\log t$, $\log\log t$ in finer regimes); $\log b$ enters only through this boundary inversion.
Thus inequality-based calibration (Markov/Ville) provides validity but not the correct asymptotic boundary for power or Bayes-risk optimality; the latter requires the KL/moderate-deviation geometry developed in Section˜3.8 and Proposition˜5.4. For a concrete worked comparison of the two regimes in a safe-testing context, see Ram [1, Example 8.1].
Utility-adjusted evidence.
If utilities $U_{1},U_{0}$ enter the decision problem, the optimal boundary becomes
$B_{t}\;\geq\;\frac{\pi_{0}\,(L_{10}-U_{0})}{\pi_{1}\,(L_{01}-U_{1})}.$
The evidence process remains unchanged; only the execution-layer boundary shifts.
Thus, within the coherent predictive subclass, likelihood-ratio evidence is the canonical representation of sequential evidence under log-loss. Inequality-calibrated thresholds (Markov/Ville) guarantee validity for any chosen boundary, but can conflict with Bayes-risk optimality: validity alone does not select the critical region that optimizes Type I/Type II tradeoffs.
3 E-values and E-processes
Remark 3.1 (Representation versus construction—a guiding distinction).
Throughout this paper, the reader should keep two layers separate. Representation layer: under coherent prediction and log-loss, the Fubini/Bayes-risk decomposition identifies the likelihood ratio (or Bayes factor) as the canonical evidence statistic; boundary selection is governed by the moderate-deviation geometry of $\log B_{t}$ (scale $\sqrt{(\log b)/t}$). Construction layer: any nonnegative supermartingale with $\operatorname{\mathbb{E}}_{0}[E_{0}]\leq 1$ is a valid E-process; Markov/Ville calibration certifies Type I control (scale $1/b$) regardless of likelihood-ratio structure. Validity (Markov/Ville) and optimality (LR/Bayes risk) are complementary, not interchangeable.
All definitions in this section live at level (L3): they require a filtered probability space.
3.1 E-variables
Definition 3.2 (E-variable).
Let $(\Omega,\mathcal{F},\operatorname{\mathbb{P}}_{H_{0}})$ be a probability space representing the null hypothesis. A nonnegative random variable $E$ is an E-variable for $H_{0}$ if $\operatorname{\mathbb{E}}_{H_{0}}[E]\leq 1$. Its realizations are called E-values.
The fundamental inference property of E-values follows from Markov’s inequality.
Theorem 3.3 (Markov bound for E-values).
Level: (L3); no bridge required.
If $E$ is an E-variable for $H_{0}$, then for all $c>0$,
$\operatorname{\mathbb{P}}_{H_{0}}(E\geq c)\leq\frac{1}{c}.$
In particular, an E-value of $c=20$ has probability at most $1/20=5\%$ under the null.
Notation reminder. When implementing a Type I risk budget $\alpha$, set $c=1/\alpha$; we write $\operatorname{\mathbb{P}}_{0}(E\geq 1/\alpha)\leq\alpha$ throughout the operational sections (Sections˜3.7 and 3.8).
Unlike $p$-values, E-values admit natural combination rules. If $E_{1},\dots,E_{m}$ are independent E-variables for the same null, their product $\prod_{i=1}^{m}E_{i}$ is again an E-variable (by the product property of expectations under independence). If $E_{1},\dots,E_{m}$ are possibly dependent E-variables for the same null, their average $\frac{1}{m}\sum_{i=1}^{m}E_{i}$ is an E-variable (by linearity of expectation). These properties make E-values a flexible currency for combining evidence across experiments and analyses [42]. Recent work has emphasized E-values as a replacement for classical $p$-values in sequential hypothesis testing, highlighting their supermartingale structure and optional-stopping validity [22].
3.2 E-processes and optional stopping
Definition 3.4 (E-process).
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\operatorname{\mathbb{P}}_{H_{0}})$ be a filtered probability space. A process $(E_{t})_{t\geq 0}$ adapted to $(\mathcal{F}_{t})$ is an E-process for $H_{0}$ if it is a nonnegative supermartingale under $H_{0}$ with $\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$.
Unless otherwise stated, we use the canonical filtration $\mathcal{F}_{t}:=\sigma(X_{1},\dots,X_{t})$ throughout (see Section˜1.6).
The central result ensuring the sequential validity of E-processes is Ville’s inequality, which extends Doob’s maximal inequality for nonnegative supermartingales [40, 14].
Theorem 3.5 (Ville’s inequality).
Level: (L3); no bridge required; requires filtration $(\mathcal{F}_{t})$ and stopping time.
If $(E_{t})_{t\geq 0}$ is an E-process under $H_{0}$, then for any $c>0$,
$\operatorname{\mathbb{P}}_{H_{0}}\Big(\sup_{t\geq 0}E_{t}\geq c\Big)\leq\frac{1}{c}.$
Consequently, for any stopping time $\tau$ (with respect to the filtration $(\mathcal{F}_{t})$),
$\operatorname{\mathbb{P}}_{H_{0}}(E_{\tau}\geq c)\leq\frac{1}{c}.$
Proof.
By the optional stopping theorem for nonnegative supermartingales, $\operatorname{\mathbb{E}}_{H_{0}}[E_{\tau\wedge T}]\leq\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$ for any bounded stopping time $\tau\wedge T$. Markov’s inequality gives $\operatorname{\mathbb{P}}_{H_{0}}(E_{\tau\wedge T}\geq c)\leq 1/c$. Letting $T\to\infty$ and applying monotone convergence yields the result for the supremum. ∎
Ville’s inequality is the foundation for anytime-valid inference: a sequential test that rejects $H_{0}$ at the first time $E_{t}\geq 1/\alpha$ controls Type I error at level $\alpha$ under arbitrary optional stopping, without any pre-specification of the sample size or stopping rule. This supermartingale perspective underlies recent proposals to reformulate hypothesis testing in terms of E-values rather than fixed significance levels [22].
Validity is not optimality.
Markov’s inequality yields $\operatorname{\mathbb{P}}_{0}(E\geq c)\leq 1/c$, hence the anytime-valid rejection rule $\tau=\inf\{t:E_{t}\geq 1/\alpha\}$. Markov/Ville inequalities turn an E-process into an anytime-valid rejection rule, guaranteeing Type I control, but they do not by themselves select a Bayes-optimal rejection region. Optimality enters through the decision problem (priors and losses), in which the likelihood ratio is the canonical statistic and the cutoff is $c^{\star}=\pi_{0}L_{10}/(\pi_{1}L_{01})$; see Section˜3.7. The rule “reject when $E_{t}\geq 1/\alpha$” is the risk-budget special case of this decision principle.
3.3 Canonical construction: likelihood ratio E-process
Likelihood-ratio processes are the canonical model-based E-processes. Under simple hypotheses they are martingales; under composite hypotheses they arise from prior-predictive mixtures as in Theorem˜2.6.
Proposition 3.6 (Likelihood ratio is an E-process under the null).
[…]
Let $H_{0}$ specify a predictive kernel $p_{0}(x_{t}\,|\,x^{t-1})$ and let $q$ be any alternative predictive kernel. Define the likelihood ratio process
$E_{t}:=\prod_{s=1}^{t}\frac{q(X_{s}\,|\,X^{s-1})}{p_{0}(X_{s}\,|\,X^{s-1})}.$
Then $(E_{t})_{t\geq 0}$ (with $E_{0}:=1$) is a nonnegative martingale under $H_{0}$, hence an E-process.
Proof.
Write $e_{s}:=q(X_{s}\,|\,X^{s-1})/p_{0}(X_{s}\,|\,X^{s-1})$ for each $s\geq 1$ and $E_{t}=\prod_{s=1}^{t}e_{s}$. Under $H_{0}$,
$\operatorname{\mathbb{E}}_{H_{0}}[e_{s}\,|\,\mathcal{F}_{s-1}]=\sum_{x_{s}}\frac{q(x_{s}\,|\,X^{s-1})}{p_{0}(x_{s}\,|\,X^{s-1})}\cdot p_{0}(x_{s}\,|\,X^{s-1})\\ =\sum_{x_{s}}q(x_{s}\,|\,X^{s-1})=1.$
Therefore $\operatorname{\mathbb{E}}_{H_{0}}[E_{t}\,|\,\mathcal{F}_{t-1}]=E_{t-1}\cdot\operatorname{\mathbb{E}}_{H_{0}}[e_{t}\,|\,\mathcal{F}_{t-1}]=E_{t-1}$, confirming the martingale property. ∎
Remark 3.7 (Additive form and the loss bridge).
On the log scale,
$\log E_{t}=\sum_{s=1}^{t}\log\left(\frac{q(X_{s}\,|\,X^{s-1})}{p_{0}(X_{s}\,|\,X^{s-1})}\right)=\ell_{p_{0}}(X^{t})-\ell_{q}(X^{t}),$
the cumulative log-loss improvement of $q$ relative to $p_{0}$. This is precisely Good’s weight of evidence (Definition˜2.5) when the two predictive specifications are $p_{0}$ and $q$.
Remark 3.8 (Girsanov change of measure).
The likelihood-ratio process $B_{t}$ is the Radon–Nikodym derivative of $P_{1}$ with respect to $P_{0}$ on $\mathcal{F}_{t}$. Sequential testing therefore monitors the density process that transports $P_{0}$ to $P_{1}$, i.e., a discrete-time Girsanov change of measure.
Representation versus construction.
The condition $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ (and its process analogue) is a validity constraint: it certifies anytime Type I control via Markov/Ville. A distinct issue is representation: when evidence is required to arise from a coherent predictive law under log-loss, the additive evidence is a log likelihood ratio (Section˜2.3), and the canonical E-process is the corresponding likelihood-ratio martingale. Validity-only constructions that satisfy $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ but do not arise from predictive kernels can yield genuine E-processes—in particular, for composite nulls where no single likelihood ratio suffices [1, 33]—but they do not inherit the Bayes-risk/weight-of-evidence semantics that the log-loss bridge provides. Cutoffs, in turn, belong to the decision layer (Section˜3.7), not the evidence layer.
3.4 Mixture E-processes: the Bayesian construction
When the alternative is composite, the Bayesian approach constructs an E-process by mixing over alternative predictive kernels.
Proposition 3.9 (Bayes factor is an E-variable under $M_{0}$).
[…]
Let $M_{0},M_{1}$ be two models with marginal likelihoods
$P_{0}(x^{n})$ $=\int L(\theta;x^{n})\,\pi_{0}(d\theta),$
$P_{1}(x^{n})$ $=\int L(\theta;x^{n})\,\pi_{1}(d\theta).$
Then the Bayes factor
$\text{BF}_{10}(X^{n}):=\frac{P_{1}(X^{n})}{P_{0}(X^{n})}$
satisfies $\operatorname{\mathbb{E}}_{M_{0}}[\text{BF}_{10}]=1$, and is therefore an E-variable under the prior-predictive (mixture) null model $M_{0}$.
Proof.
Under $M_{0}$, the marginal distribution of $X^{n}$ is $P_{0}$. Therefore
$\operatorname{\mathbb{E}}_{M_{0}}[\text{BF}_{10}]=\int\frac{P_{1}(x^{n})}{P_{0}(x^{n})}\,P_{0}(dx^{n})=\int P_{1}(x^{n})\,dx^{n}=1.$
This establishes validity under the model-averaged null $M_{0}$; for composite null classes $\mathcal{P}$, uniform validity $\sup_{P\in\mathcal{P}}\operatorname{\mathbb{E}}_{P}[E_{t}]\leq 1$ is a strictly stronger requirement and generally distinct from mixture-null validity. Hence sequential Bayes factors inherit anytime validity, multiplicative combination, and Ville’s inequality.
Composite nulls: mixture validity vs uniform validity.
When the null is specified as a set of measures $\mathcal{P}$ (a composite hypothesis), the e-value literature typically requires uniform validity, $\operatorname{\mathbb{E}}_{P}[E_{t}]\leq 1$ for every $P\in\mathcal{P}$, which is strictly stronger than validity under a prior-predictive mixture null $M_{0}=\int P_{\theta}\,\pi_{0}(d\theta)$ (the setting of Proposition˜3.9). In this uniform regime, generic Bayes factors need not be e-variables for $\mathcal{P}$. Nevertheless, a likelihood-ratio representation remains canonical in an appropriate minimax sense: Larsson et al. [27] show that for arbitrary $\mathcal{P}$ and fixed alternative $Q$, there exists a log-optimal (“numeraire”) e-variable $X^{*}$ that can be represented as a likelihood ratio $dQ/dP^{*}$ for a distinguished representative $P^{*}$ (a generalized reverse information projection/bipolar construction). This complements our mixture-null (Bayes-risk) representation: both regimes yield LR-type canonical objects, but with different null semantics. Accordingly, statements of the form “BF == e-value” are correct under the mixture (prior-predictive) null semantics, and generally false under uniform composite-null semantics unless one uses a numeraire/projection construction (or very special priors).
By Proposition˜2.3, the log Bayes factor decomposes as
$\log\text{BF}_{10}=\ell_{P_{0}}(x^{n})-\ell_{P_{1}}(x^{n}),$
the difference of two negative log-marginal likelihoods, each of which is a log-sum-exp aggregation over the respective parameter space.
Remark 3.10 (LR martingale symmetry).
Let $B_{t}:=\frac{dP_{1}}{dP_{0}}\big|_{\mathcal{F}_{t}}$ (e.g., a Bayes factor or likelihood ratio against a fixed or mixture alternative). Then $(B_{t})_{t\geq 0}$ is a nonnegative $P_{0}$-martingale with $\operatorname{\mathbb{E}}_{0}[B_{t}]=1$ for all $t$. Moreover, $(1/B_{t})_{t\geq 0}$ is a nonnegative $P_{1}$-martingale with $\operatorname{\mathbb{E}}_{1}[1/B_{t}]=1$ for all $t$. Thus the likelihood ratio is a two-sided evidence currency: it is simultaneously an E-process for $H_{0}$ and, after inversion, an E-process for $H_{1}$. [^1] In finance, the same LR/martingale viewpoint underlies empirical work on pricing kernels and martingale restrictions; see, e.g., Bossaerts [5] and subsequent option-kernel testing literature.
Thus each hypothesis admits its own canonical evidence martingale.
3.5 Worked example: Bernoulli coin with Beta prior
Let $X_{1},X_{2},\ldots$ be sequential observations in $\{0,1\}$, and consider the simple null $H_{0}:X_{s}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Bernoulli}(1/2)$. Let $S_{t}:=\sum_{s=1}^{t}X_{s}$. For a composite alternative $H_{1}:X_{s}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Bernoulli}(p)$ with $p$ unknown, equip $p$ with a $\text{Beta}(a,b)$ prior, $a,b>0$. The Bayes factor at time $t$ is
$E_{t}=\frac{\int_{0}^{1}p^{S_{t}}(1-p)^{t-S_{t}}\,\pi(dp)}{(1/2)^{t}}=2^{t}\,\frac{B(a+S_{t},\,b+t-S_{t})}{B(a,b)},$
where $B(u,v):=\int_{0}^{1}p^{u-1}(1-p)^{v-1}\,dp$ is the Beta function.
Under $H_{0}$, $(E_{t})_{t\geq 1}$ is a nonnegative martingale, hence an E-process, and Ville’s inequality yields
$\operatorname{\mathbb{P}}_{H_{0}}\!\left(\sup_{t\geq 1}E_{t}\geq\frac{1}{\alpha}\right)\leq\alpha.$
A sequential test that rejects $H_{0}$ at the first time $E_{t}\geq 1/\alpha$ controls Type I error at level $\alpha$ under arbitrary optional stopping; the experimenter may peek at the data at every step without inflating the error rate.
3.6 Good’s paradox for p-values; stopping-rule invariance via E-processes
Good highlighted a paradox for p-values under optional stopping: the stopping-rule dependence of classical procedures, and independence of Bayesian and E-value methods, was central in debates. By the law of the iterated logarithm, a frequentist using optional stopping can always reject a true null given patience [19].
Specifically, for a random walk $S_{n}=\sum_{i=1}^{n}X_{i}$ with $X_{i}$ i.i.d. having mean zero and variance $\sigma^{2}$,
$\limsup_{n\to\infty}\frac{S_{n}}{\sigma\sqrt{2n\log\log n}}=1\quad\text{almost surely},$
so the standardized sum $S_{n}/(\sigma\sqrt{n})$ will eventually exceed any fixed threshold with probability one, even under the null.
Good’s point was that Bayesian inference, based on the likelihood principle, is immune to this manipulation: the Bayes factor depends only on the data observed, not on the stopping rule.
E-values clarify this tension by providing frequentist error control that is valid under arbitrary stopping while also coinciding with Bayesian evidence. Since Bayes factors are E-values under the mixture (prior-predictive) null (Proposition˜3.9), and E-processes maintain Type I error control under optional stopping via Ville’s inequality (Theorem˜3.5), the E-value framework shows that one can have both: evidential measures that are likelihood-based and sequential validity that does not require pre-specification of stopping rules. The “safe testing” programme formalizes this perspective, providing E-value constructions with explicit robustness guarantees [21, 39].
In particular, Good’s example highlights how fixed-horizon $p$-value calibration can break under repeated peeking unless adjusted. E-processes, by contrast, satisfy Ville’s inequality under arbitrary stopping, so Type I error control is preserved regardless of the stopping rule. The remaining question for E-values is not stopping validity but boundary selection: given that validity is guaranteed, which threshold yields Bayes-risk-optimal decisions? Our distinction in the sequel concerns precisely this optimality layer: Markov/Ville certify validity for a chosen boundary, but do not by themselves determine Bayes-risk-aligned rejection regions when LR structure is available.
3.7 Decision-theoretic cutoffs
Consider binary action $\delta_{t}\in\{0,1\}$ at time $t$ with losses $L_{10}$ (reject $H_{0}$ when $H_{0}$ true) and $L_{01}$ (fail to reject when $H_{1}$ true), and prior odds $\pi_{1}/\pi_{0}$. Bayes rule rejects $H_{0}$ when posterior odds exceed the loss ratio:
$\frac{\pi_{1}}{\pi_{0}}\,B_{t}\;\geq\;\frac{L_{10}}{L_{01}}\quad\Longleftrightarrow\quad B_{t}\;\geq\;c^{\star}:=\frac{\pi_{0}\,L_{10}}{\pi_{1}\,L_{01}}.$
The cutoff is decision-theoretic: once $(\pi_{0},\pi_{1},L_{10},L_{01})$ are fixed, the threshold $c^{\star}$ is fixed; $t$ affects only the sampling distribution of $B_{t}$.
The anytime-valid testing rule “reject at the first time $B_{t}\geq 1/\alpha$” is the special case corresponding to a risk-budget choice that sets $c^{\star}=1/\alpha$. This is the operational form of the same decision principle when one requires uniform Type I control under optional stopping via Ville’s inequality. Cutoffs belong to the decision layer; the representation $E_{t}=\Lambda_{t}$ (Theorem˜2.6) belongs to the evidence layer.
3.8 Critical region geometry for E-values
The critical region for sequential evidence is governed by two distinct asymptotic geometries. Failing to distinguish them is the source of the efficiency gap noted in Remark˜2.7.
Validity geometry Optimality geometry
(Markov/Ville) (Likelihood ratio / mod. dev.)
Acts on $E_{t}$ directly $\log E_{t}$
Tail scale Polynomial: $\leq 1/b$ Exponential: $\sqrt{(\log b)/t}$
Determines Type I certificate Type I/II frontier (power)
Limitation Conservative; not Bayes-optimal Requires LR structure
Here $b$ is a fixed evidence cutoff (e.g. $b=1/\alpha$); all asymptotics are in the sample size $t$. The $\log b$ term arises solely from the boundary inversion $B_{t}\geq b\Leftrightarrow\log B_{t}\geq\log b$ and is not an independent asymptotic variable.
Geometry 1: Markov/Ville (validity calibration).
For any nonnegative E-process,
$\operatorname{\mathbb{P}}_{0}\!\left(\sup_{s\leq t}E_{s}\geq b\right)\leq\frac{1}{b}.$
Markov/Ville certify validity for a chosen cutoff; setting $b=1/\alpha$ gives the anytime-valid certificate $\operatorname{\mathbb{P}}_{0}(\sup E_{s}\geq 1/\alpha)\leq\alpha$. This bound operates on $E_{t}$ at a polynomial $\sim\!1/b$ scale. It guarantees Type I control but is agnostic about likelihood-ratio structure and therefore does not optimize the critical region.
Geometry 2: Likelihood ratio / moderate deviations (optimality calibration).
When $E_{t}=B_{t}$ is a likelihood ratio,
$\frac{1}{t}\log B_{t}\;\to\;D_{\operatorname{KL}}(P_{1}\|P_{0})\quad\text{a.s.\ under }P_{1},$
and moderate-deviation theory governs the fluctuations: deviations of $(1/t)\log B_{t}$ from its KL slope are of order $\sqrt{(\log b)/t}$—an exponentially sharper scale than the polynomial $1/b$ of Markov/Ville. The connection is direct: crossing $B_{t}\geq b$ is equivalent to $\log B_{t}\geq\log b$, and since $(1/t)\log B_{t}\approx D_{\operatorname{KL}}(P_{1}\|P_{0})$ under $P_{1}$, the threshold $\log b$ is naturally calibrated against $t\,D_{\operatorname{KL}}$. This is the appropriate scaling for boundary selection and power analysis. [^2] A concrete comparison: for the fixed threshold $b=20$ at $t=100$ observations, the Markov/Ville bound gives $\operatorname{\mathbb{P}}_{0}(E_{t}\geq 20)\leq 1/20=0.05$, while the moderate-deviation fluctuation scale is $\sqrt{(\log 20)/100}\approx 0.17$ nats—the two operate on entirely different objects ($E_{t}$ versus $(1/t)\log B_{t}$) and at different scales. See also Proposition˜5.4.
Remark 3.11 (Two geometries, one evidence process).
The Markov/Ville boundary is polynomial and universally valid but may be loose; the likelihood-ratio boundary is governed by moderate-deviation theory and yields exponentially sharper boundaries. Procedures relying solely on Markov/Ville calibration may therefore be potentially loose in finite samples, since they do not exploit the KL growth structure available for likelihood-ratio processes. The evidence process $B_{t}$ is the same in both cases; what differs is the geometry used to select and evaluate the critical region.
In short: Ville/Markov tells you that a boundary $b=1/\alpha$ is valid; it does not tell you it is optimal. Optimal boundaries track $\log B_{t}$ and therefore inherit $\sqrt{(\log b)/t}$-scale moderate deviations around the KL slope.
Normalization.
Rate statements are naturally expressed on the log scale: we study $(1/t)\log E_{t}$ (or $(1/t)\log B_{t}$) rather than $E_{t}/t$. This places the Markov/Ville calibration (on $E_{t}$) and KL growth (on $\log E_{t}$) in a common time-normalized comparison.
Why the Markov region is not an evidential region.
The inequality $\operatorname{\mathbb{P}}_{0}(\sup_{s\leq t}E_{s}\geq 1/\alpha)\leq\alpha$ is a validity certificate for an execution rule; it does not convert $E_{t}$ into a tail probability nor identify a Bayes-optimal rejection region. Interpreting $E_{t}$ as “1/p1/p” conflates two typed objects: Bayes factors compare models relatively via likelihood ratios, whereas Markov/Ville bounds are one-sided guarantees that ignore the alternative and therefore cannot optimize Type II error along the frontier.
Optimal threshold scaling.
The $\sqrt{(\log b)/t}$ moderate-deviation scale is not merely descriptive: in related Bayes-risk calibrations one can solve explicitly for the optimizing critical level, yielding thresholds of order $\sqrt{(\lambda+m+1)\log n}$ rather than $1/n$-type calibrations, where $\lambda,m$ are problem-specific parameters; see Datta et al. [10] for a worked-out instance.
3.9 Validity vs. efficiency: boundary design under LR structure
E-process thresholds such as $E_{\tau}\geq 1/\alpha$ arise from Ville’s inequality and provide exact Type I control under arbitrary stopping. This is a validity-layer guarantee: it does not depend on distributional assumptions beyond the e-process property.
However, when the evidence process admits a likelihood-ratio representation (simple LR, mixture Bayes factor, or numeraire LR against a representative $P^{*}$), the distribution of $\log E_{t}$ under alternatives exhibits a large-deviation and moderate-deviation structure. In that regime, boundary selection becomes an efficiency problem rather than a validity problem.
Specifically, if $\log E_{t}$ satisfies a moderate deviation principle with rate governed by the KL slope, then calibrating thresholds solely via the universal $1/\alpha$ scale may sacrifice efficiency relative to boundaries tuned to the moderate-deviation geometry.
The distinction is therefore not between correct and incorrect thresholds, but between universal validity guarantees and asymptotically efficient boundary design when LR structure is available.
When $E_{t}$ has LR-type structure (simple LR, mixture BF, or numeraire LR against a representative $P^{*}$), the boundary problem reduces to calibrating deviations of $\log E_{t}$, which inherit the KL geometry of the underlying model. Thus the appropriate scaling is dictated by moderate-deviation behavior rather than solely by Markov bounds.
This observation does not contradict the validity of Ville-calibrated e-processes; rather, it identifies conditions under which sharper boundary design is possible.
Validity Efficiency
Ville/Markov Exact Possibly loose
MDP/KL calibration Exact (when e-process) KL-optimal under LR structure
3.10 Sequential monitoring applications
In sequential monitoring (e.g., clinical trials with interim analyses), Ville’s inequality permits continuous monitoring without $\alpha$-spending boundaries [43, 25]:
$\operatorname{\mathbb{P}}\!\left(\sup_{n\geq 1}E_{n}\geq\frac{1}{\alpha}\right)\leq\alpha.$
Adaptive designs are accommodated provided that all design modifications are predictable with respect to the filtration, so that the supermartingale property is preserved [6].
3.11 Comparison with $p$-values
The comparison between E-values and fixed-horizon $p$-values can be organized around three axes [42, 1].
Optional stopping.
If $(E_{n})$ is an E-process, Ville’s inequality provides valid Type I error control at arbitrary stopping times. Fixed-horizon $p$-values generally require adjustments (e.g., $\alpha$-spending) for sequential use.
Combination.
Independent E-variables can be multiplied; possibly dependent E-variables can be averaged. Both operations preserve the E-variable property. Combining $p$-values requires distributional assumptions or conservative bounds (e.g., Bonferroni, Fisher’s method).
Relation to Bayes factors and safe tests.
Bayes factors are E-variables under the mixture (prior-predictive) null (Proposition˜3.9), inheriting all sequential validity properties. The safe testing framework studies E-value constructions with robustness guarantees under optional stopping and model misspecification [21].
3.12 Duality with confidence sequences
E-processes are dual to confidence sequences, sequences of confidence sets $(C_{n})_{n\geq 1}$ that are simultaneously valid at all sample sizes: $\operatorname{\mathbb{P}}(\theta\in C_{n}\;\text{for all }n\geq 1)\geq 1-\alpha$ [23]. The duality is constructive: from any E-process $(E_{n}(\theta))_{n\geq 1}$ indexed by a parameter $\theta$, one obtains a $(1-\alpha)$-confidence sequence by inverting $C_{n}:=\{\theta:E_{n}(\theta)<1/\alpha\}$; conversely, any confidence sequence induces a family of E-processes via the same inversion [1]. The same supermartingale machinery underlies both sequential tests and sequential estimation [44].
Bayesian perspective.
From a Bayesian perspective, Bayes factors arise as likelihood-ratio E-processes after marginalization over parameters. E-values extend this viewpoint by retaining anytime-valid error control under optional stopping while remaining interpretable as sequential likelihood ratios or mixtures thereof, provided the construction is predictably normalized.
Our point in the following sections is not to oppose e-values or the broader e-process programme, but to separate (i) validity certificates (Markov/Ville) from (ii) representational semantics (likelihood-ratio structure under coherent prediction) and (iii) decision-layer boundary choice (Bayes-risk-optimal cutoffs). We view these as complementary, not competing, layers of sequential inference.
4 Composition, amalgamation, and obstructions in the evidence class
This section formalizes the algebraic structure of the evidence class: which operations on E-processes preserve validity, and which provably do not.
4.1 The evidence class and its basic algebra
Definition 4.1 (Evidence class under $H_{0}$).
Let $H_{0}$ denote the null model class (as in §3, Definition˜3.2). Define the evidence class
$\mathcal{E}(H_{0}):=\{E\geq 0:\operatorname{\mathbb{E}}_{H_{0}}[E]\leq 1\}.$
Similarly, on a filtered space $(\Omega,\mathcal{F},(\mathcal{F}_{t}),\operatorname{\mathbb{P}})$, define the process evidence class
$\mathcal{E}^{\mathrm{proc}}(H_{0}):=\{(E_{t})_{t\geq 0}:E_{t}\geq 0,\;(E_{t})\ \text{is a supermartingale under }H_{0},\;\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1\}.$
Remark 4.2 (Terminology: class, not cone).
The set $\mathcal{E}(H_{0})$ is convex but not a cone in the strict algebraic sense, since the closure $cE\in\mathcal{E}(H_{0})$ holds only for scalars $c\in(0,1]$; scaling by $c>1$ may violate $\operatorname{\mathbb{E}}_{H_{0}}[cE]\leq 1$. We therefore use “class” rather than “cone.”
Proposition 4.3 (Convex closure (amalgamation)).
  1. (i)
    If $E_{1},\dots,E_{m}\in\mathcal{E}(H_{0})$ and $w_{i}\geq 0$ with $\sum_{i=1}^{m}w_{i}=1$, then $E:=\sum_{i=1}^{m}w_{i}E_{i}\in\mathcal{E}(H_{0})$.
  2. (ii)
    If $(E^{i}_{t})\in\mathcal{E}^{\mathrm{proc}}(H_{0})$ for $i=1,\dots,m$, then $E_{t}:=\sum_{i=1}^{m}w_{i}E^{i}_{t}$ defines $(E_{t})\in\mathcal{E}^{\mathrm{proc}}(H_{0})$.
Proof.
For (i), linearity of expectation gives $\operatorname{\mathbb{E}}_{H_{0}}[E]=\sum_{i=1}^{m}w_{i}\operatorname{\mathbb{E}}_{H_{0}}[E_{i}]\leq\sum_{i=1}^{m}w_{i}=1$, and $E\geq 0$ since each $E_{i}\geq 0$ and $w_{i}\geq 0$. For (ii), $E_{t}\geq 0$ is immediate, and the conditional expectation satisfies
$\operatorname{\mathbb{E}}_{H_{0}}[E_{t}\,|\,\mathcal{F}_{t-1}]=\sum_{i=1}^{m}w_{i}\operatorname{\mathbb{E}}_{H_{0}}[E^{i}_{t}\,|\,\mathcal{F}_{t-1}]\leq\sum_{i=1}^{m}w_{i}E^{i}_{t-1}=E_{t-1},$
confirming the supermartingale property. The initial condition $\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$ follows as in (i). ∎
Remark 4.4 (Relation to “merging” results).
Proposition˜4.3 shows that validity of averaging and mixtures is a closure property of the evidence class. Recent work on combining and merging E-values [42, 1] studies finer admissibility questions, which merging functions are optimal within the class, but the basic algebra captured here underlies all such results.
4.2 Anytime validity as a compositional consequence of Ville
Theorem 4.5 (Anytime validity under convex combinations).
Let $(E_{t})\in\mathcal{E}^{\mathrm{proc}}(H_{0})$. Then for all $c>0$,
$\operatorname{\mathbb{P}}_{H_{0}}\Big(\sup_{t\geq 0}E_{t}\geq c\Big)\leq\frac{1}{c}.$
Moreover, for any stopping time $\tau$, $E_{\tau}\in\mathcal{E}(H_{0})$ and $\operatorname{\mathbb{P}}_{H_{0}}(E_{\tau}\geq c)\leq 1/c$.
Proof.
This is Theorem˜3.5 applied to $(E_{t})$. The stopped variable $E_{\tau}$ satisfies $\operatorname{\mathbb{E}}_{H_{0}}[E_{\tau}]\leq\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$ by the optional stopping theorem for nonnegative supermartingales, confirming $E_{\tau}\in\mathcal{E}(H_{0})$. ∎
Remark 4.6 (What is “anytime validity”?).
What is called “anytime-valid” or “always-valid” inference is the statement of Theorem˜4.5: it follows from nonnegative supermartingale structure and Ville’s inequality (Theorem˜3.5). In particular, likelihood-ratio and Bayes-factor processes are anytime-valid by construction (Propositions˜3.6 and 3.9), and convex mixtures preserve this property (Proposition˜4.3).
4.3 Sequential composition (stitching) of evidence processes
Arbitrary products of supermartingales need not be supermartingales; in particular, the product of two independent E-processes is not generally an E-process without additional conditions. Independence of terminal values is not the same as conditional independence with respect to a shared filtration; without joint adaptedness the supermartingale property can fail. The safe sequential operation is stitching: restarting a new E-process after a data-dependent stopping time.
Definition 4.7 (Stitched evidence process).
Let $\tau$ be a stopping time. Suppose $(E^{(1)}_{t})_{t\geq 0}\in\mathcal{E}^{\mathrm{proc}}(H_{0})$ and, conditionally on $\mathcal{F}_{\tau}$, we define a second process $(E^{(2)}_{t})_{t\geq 0}$ such that $(E^{(2)}_{t})$ is an E-process on the shifted filtration $(\mathcal{F}_{\tau+t})_{t\geq 0}$ under $H_{0}$ with $E^{(2)}_{0}=1$. Define the stitched process
$\tilde{E}_{t}:=\begin{cases}E^{(1)}_{t},&t\leq\tau,\\ E^{(1)}_{\tau}\cdot E^{(2)}_{t-\tau},&t>\tau.\end{cases}$
Proposition 4.8 (Stitching validity).
Under Definition˜4.7, $(\tilde{E}_{t})_{t\geq 0}\in\mathcal{E}^{\mathrm{proc}}(H_{0})$. Thus $\tilde{E}_{\tau}$ and $\sup_{t}\tilde{E}_{t}$ satisfy the Ville bounds of Theorem˜4.5.
Proof.
For $t\leq\tau$, $\tilde{E}_{t}=E^{(1)}_{t}$ is a supermartingale by assumption. For $t>\tau$, write $\tilde{E}_{t}=E^{(1)}_{\tau}\cdot E^{(2)}_{t-\tau}$. Taking conditional expectations on the shifted filtration:
$\operatorname{\mathbb{E}}_{H_{0}}[\tilde{E}_{t}\,|\,\mathcal{F}_{t-1}]=E^{(1)}_{\tau}\cdot\operatorname{\mathbb{E}}_{H_{0}}[E^{(2)}_{t-\tau}\,|\,\mathcal{F}_{t-1}]\leq E^{(1)}_{\tau}\cdot E^{(2)}_{t-1-\tau}=\tilde{E}_{t-1},$
where the inequality uses the supermartingale property of $(E^{(2)}_{s})$ on the shifted filtration. The boundary at $t=\tau$ is handled by $E^{(2)}_{0}=1$, giving $\tilde{E}_{\tau}=E^{(1)}_{\tau}\cdot 1$. ∎
4.4 Gibbs/Bayes amalgamation as a canonical constructor
Proposition 4.9 (Mixture of likelihood-ratio E-processes).
Let $\{q_{\theta}(\cdot\mid\cdot)\}_{\theta\in\Theta}$ be a family of alternative predictive kernels and let $p_{0}(\cdot\mid\cdot)$ be the null kernel. For each $\theta$, define the likelihood-ratio process
$E^{\theta}_{n}:=\prod_{t=1}^{n}\frac{q_{\theta}(X_{t}\mid X^{t-1})}{p_{0}(X_{t}\mid X^{t-1})},$
which is an E-process under $H_{0}$ (Proposition˜3.6). For any prior $\pi$ on $\Theta$, the mixture
$E_{n}:=\int_{\Theta}E^{\theta}_{n}\,\pi(d\theta)$
is also an E-process under $H_{0}$.
Proof.
By Fubini’s theorem and the martingale property of each $E^{\theta}_{n}$:
$\operatorname{\mathbb{E}}_{H_{0}}[E_{n}\,|\,\mathcal{F}_{n-1}]=\int_{\Theta}\operatorname{\mathbb{E}}_{H_{0}}[E^{\theta}_{n}\,|\,\mathcal{F}_{n-1}]\,\pi(d\theta)\leq\int_{\Theta}E^{\theta}_{n-1}\,\pi(d\theta)=E_{n-1}.$
Nonnegativity and the initial condition $E_{0}=1$ are immediate. ∎
Remark 4.10 (Connection to the monoidal bridge).
Proposition˜4.9 is the process-level counterpart of the mixture identity in Proposition˜2.3. On the log scale, Bayesian marginalization corresponds to a log-sum-exp aggregation of losses, yielding a coherent amalgamation rule that respects both mixtures (L2) and evidence processes (L3). These closure properties are the formal content underlying the bridge summary in Table˜2 and Figure˜1.
4.5 Non-compositions and typed obstructions
Remark 4.11 (Operations that do not preserve E-validity).
The following operations generally fail to preserve membership in $\mathcal{E}(H_{0})$ or $\mathcal{E}^{\mathrm{proc}}(H_{0})$ unless supplemented by additional calibration:
  1. (i)
    $\min_{i}E_{i}$ and $\max_{i}E_{i}$: pointwise max/min are nonlinear operations that violate convex closure of the evidence class and need not preserve Ville’s inequality;
  2. (ii)
    thresholding $E\mapsto\mathbf{1}\{E\geq c\}$ or truncations not designed as calibrators;
  3. (iii)
    naive averaging or optional-stopping of p-values (lack of supermartingale structure);
  4. (iv)
    composing set-valued certificates (e.g., conformal prediction sets) as if they were predictive states (failure of descent under refinement/stopping).
These failures are not pathologies but reflect typed mismatch: the objects produced by these operations do not live in the evidence class, and no restriction or gluing maps exist that would make them stable under refinement and stopping. The non-bridges catalogued in §8.3 are specific instances of this general obstruction principle.
4.6 Second-moment refinement: variance-aware evidence
Definition 4.12 ($L^{2}$-evidence class).
For $C\in(1,\infty]$, define
$\mathcal{E}_{2}(H_{0};C):=\{E\geq 0:\operatorname{\mathbb{E}}_{H_{0}}[E]\leq 1,\ \operatorname{\mathbb{E}}_{H_{0}}[E^{2}]\leq C\}.$
For processes, one may analogously define an $L^{2}$-refined class by requiring $\operatorname{\mathbb{E}}_{H_{0}}[(E_{t})^{2}]\leq C$ at each time $t$, or by controlling the predictable quadratic variation $\langle\log E\rangle_{t}$.
Proposition 4.13 (Convex closure preserves second moments).
If $E_{i}\in\mathcal{E}_{2}(H_{0};C_{i})$ and $E=\sum_{i}w_{i}E_{i}$ with $w_{i}\geq 0$, $\sum_{i}w_{i}=1$, then $E\in\mathcal{E}_{2}(H_{0};\sum_{i}w_{i}C_{i})$.
Proof.
The expectation bound follows from Proposition˜4.3. For second moment, Jensen’s inequality with $x\mapsto x^{2}$ gives
$\operatorname{\mathbb{E}}_{H_{0}}[E^{2}]=\operatorname{\mathbb{E}}_{H_{0}}\Big[\Big(\sum_{i}w_{i}E_{i}\Big)^{2}\Big]\leq\sum_{i}w_{i}\operatorname{\mathbb{E}}_{H_{0}}[E_{i}^{2}]\leq\sum_{i}w_{i}C_{i},$
where the first inequality uses $(\sum w_{i}a_{i})^{2}\leq\sum w_{i}a_{i}^{2}$ for $w_{i}\geq 0$, $\sum w_{i}=1$. ∎
Corollary 4.14 (Convex mixtures preserve $L^{2}$-control).
If $E_{i}\in\mathcal{E}_{2}(H_{0};C)$ for all $i$ and $E=\sum_{i}w_{i}E_{i}$ with $w_{i}\geq 0$, $\sum w_{i}=1$, then $E\in\mathcal{E}_{2}(H_{0};C)$.
Proof.
Immediate from Proposition˜4.13 with $C_{i}=C$ for all $i$: $\sum_{i}w_{i}C_{i}=C\sum_{i}w_{i}=C$. ∎
Thus amalgamation within the variance-controlled class yields tighter deviation bounds (Bernstein-type rather than Markov-type) and sharper confidence sequences—all without leaving $L^{2}$.
Remark 4.15 (Why variance control is a refinement).
Second-moment control yields sharper tail bounds and tighter confidence sequences [23, 44], but requires assumptions beyond the evidence class. This is a refinement, not a replacement.
The preceding section identified closure properties of evidence objects. We now turn to the geometric invariant that governs their rates: the KL divergence.
5 KL geometry and information-regularized composition
KL divergence governs three phenomena: exponential rarity of atypical empirical distributions (Sanov), growth rates of likelihood-based evidence, and the variational structure of Bayesian updating.
5.1 Empirical laws and variational structure
Let $X^{n}=(X_{1},\dots,X_{n})$ be realizations (Level L1), and let
$\hat{P}_{n}:=\frac{1}{n}\sum_{t=1}^{n}\delta_{X_{t}}$
denote the empirical measure. Let $\mathcal{P}(\mathcal{X})$ denote the model space (Level L2), and let $\Pi_{0}$ be a prior on $\mathcal{P}(\mathcal{X})$.
Define the Gibbs posterior
$\Pi_{n}(dP)\propto\exp\{-nD_{\operatorname{KL}}(\hat{P}_{n}\|P)\}\,\Pi_{0}(dP).$ (5.1)
Proposition 5.1 (Variational characterization).
[…]
The posterior $\Pi_{n}$ in (5.1) minimizes the functional
$Q\mapsto\operatorname{\mathbb{E}}_{Q}\!\big[D_{\operatorname{KL}}(\hat{P}_{n}\|P)\big]+\frac{1}{n}D_{\operatorname{KL}}(Q\|\Pi_{0})$
over probability measures $Q$ on $\mathcal{P}(\mathcal{X})$.
Proof.
This is the Gibbs variational principle (a consequence of Donsker–Varadhan duality [13]). For any measurable $f$ and reference measure $\mu$, the minimizer of $Q\mapsto\operatorname{\mathbb{E}}_{Q}[f]+D_{\operatorname{KL}}(Q\|\mu)$ is $Q^{\star}(dP)\propto e^{-f(P)}\mu(dP)$, achieved by completing the square in relative entropy. Setting $f(P)=nD_{\operatorname{KL}}(\hat{P}_{n}\|P)$ and $\mu=\Pi_{0}$ yields exactly $\Pi_{n}$. ∎
Remark 5.2 (Connection to the monoidal bridge).
Proposition˜5.1 is the higher-level counterpart of Proposition˜2.3: Bayesian marginalization is the canonical KL-coherent amalgamation of (L1) and (L2).
5.2 Large deviations and de Finetti geometry
Under exchangeability, Sanov (Theorem˜5.3) governs rarity of empirical measures and inverse Sanov (Theorem˜5.7) governs posterior concentration. The Gibbs form (5.1) is the I-projection of the prior onto the empirical constraints; see Polson and Zantedeschi [31] for the full dual large-deviation development.
The key pair is:
[…]
Same functional, different direction of conditioning.
5.3 Sanov’s theorem: rarity of empirical distributions
Theorem 5.3 (Sanov’s theorem, finite alphabet).
Level: (L1)++(L2); bridge: empirical measure $\hat{P}_{n}$.
Let $\mathcal{X}$ be a finite alphabet, let $X_{1},X_{2},\ldots$ be i.i.d. draws from $P\in\mathcal{P}(\mathcal{X})$, and let $\hat{P}_{n}$ denote the empirical distribution. Equip $\mathcal{P}(\mathcal{X})$ with the topology induced by total variation (equivalently, the Euclidean topology on the simplex). For any Borel set $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{X})$,
$\operatorname{\mathbb{P}}_{P}(\hat{P}_{n}\in\mathcal{Q})\asymp\exp\big(-n\inf_{Q\in\mathcal{Q}}D_{\operatorname{KL}}(Q\|P)\big),$
in the precise sense that for the interior $\mathcal{Q}^{\circ}$ and closure $\overline{\mathcal{Q}}$:
$\limsup_{n\to\infty}\frac{1}{n}\log\operatorname{\mathbb{P}}_{P}(\hat{P}_{n}\in\overline{\mathcal{Q}})$ $\leq-\inf_{Q\in\overline{\mathcal{Q}}}D_{\operatorname{KL}}(Q\|P),$ (5.2)
$\liminf_{n\to\infty}\frac{1}{n}\log\operatorname{\mathbb{P}}_{P}(\hat{P}_{n}\in\mathcal{Q}^{\circ})$ $\geq-\inf_{Q\in\mathcal{Q}^{\circ}}D_{\operatorname{KL}}(Q\|P).$ (5.3)
The rate function is entirely determined by KL divergence, the central object of Shannon’s information theory. Sanov’s theorem says that the probability of the empirical distribution landing in a set $\mathcal{Q}$ decays exponentially in $n$, at a rate given by the KL divergence from $P$ to the closest distribution in $\mathcal{Q}$.
5.4 Evidence growth: law of large numbers for $\log E_{n}$
Proposition 5.4 (Evidence growth rate under an alternative).
[…]
Let $P_{0},P_{1}$ be distributions on a finite alphabet $\mathcal{X}$ with $P_{1}\ll P_{0}$. Consider the likelihood ratio E-process $E_{n}=\prod_{i=1}^{n}P_{1}(X_{i})/P_{0}(X_{i})$ where $X_{1},X_{2},\ldots$ are i.i.d. under $P_{1}$. Then
$\frac{1}{n}\log E_{n}\xrightarrow{\text{a.s.}}D_{\operatorname{KL}}(P_{1}\|P_{0})\quad\text{under }P_{1}.$
Proof.
By the strong law of large numbers, $\frac{1}{n}\log E_{n}=\frac{1}{n}\sum_{i=1}^{n}\log\frac{P_{1}(X_{i})}{P_{0}(X_{i})}\to\operatorname{\mathbb{E}}_{P_{1}}\!\left[\log\frac{P_{1}(X)}{P_{0}(X)}\right]=D_{\operatorname{KL}}(P_{1}\|P_{0})$ almost surely. ∎
Thus $D_{\operatorname{KL}}(P_{1}\|P_{0})$ is the almost-sure growth rate functional of likelihood-based E-processes. This characterization is pathwise and hence independent of stopping rules; Proposition˜B.9 extends the result to the misspecified case.
Remark 5.5 (KL growth and the moderate-deviation boundary).
The KL growth rate implies that the fluctuations of $(1/t)\log B_{t}$ around $D_{\operatorname{KL}}(P_{1}\|P_{0})$ occur on the $\sqrt{(\log b)/t}$ scale (moderate deviations), not the polynomial $1/b$ scale of Markov/Ville. This is precisely the distinction that separates Bayes-risk-optimal boundary selection (governed by KL/moderate deviations) from validity-only calibration (governed by Markov/Ville); see Remark˜2.7 in Section 2.3 and the two-geometry framework of Section˜3.8.
Rate optimality of the likelihood ratio.
The growth rate $D_{\operatorname{KL}}(P_{1}\|P_{0})$ is not only the natural growth rate but also the optimal one. No E-process for $H_{0}$ can achieve an almost-sure growth rate exceeding $D_{\operatorname{KL}}(P_{1}\|P_{0})$ under $P_{1}$ when both hypotheses are simple, because $\operatorname{\mathbb{E}}_{0}[E_{n}]\leq 1$ constrains the mean log-increment to at most $D_{\operatorname{KL}}(P_{1}\|P_{0})$ per observation; the likelihood-ratio process attains this bound with equality. Under misspecification ($P_{\mathrm{true}}\neq P_{1}$), the growth rate becomes $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})$ (Proposition˜B.9), which can vanish or turn negative when the alternative is farther from the truth than the null—a direct manifestation of the misspecification non-bridge discussed in Section˜8.
Mixtures improve tails beyond $1/c$.
Markov/Ville bounds yield polynomial-style calibration (e.g. $\operatorname{\mathbb{P}}_{0}(B_{t}\geq c)\leq 1/c$), but when $B_{t}$ is a (mixture) likelihood ratio, classical likelihood theory implies exponential concentration of $\log B_{t}$ around its KL rate (Chernoff/Wilks regimes), so that deviations scale on the order $\sqrt{\log b/t}$ rather than the polynomial $1/t$ rate of Markov-based bounds.
Remark 5.6 (Good’s die and the KL growth rate).
Good [18] illustrated the weight of evidence with a die example: testing a fair die ($P_{0}$: uniform) against a loaded alternative ($P_{1}$: biased toward one face). The accumulated weight of evidence $W(x^{n})=\sum_{t=1}^{n}\log\frac{P_{1}(x_{t})}{P_{0}(x_{t})}$ grows at rate $D_{\operatorname{KL}}(P_{1}\|P_{0})$ by the proposition above. Good’s example thus contains the seed of the modern E-process framework: the additive accumulation of log-likelihood ratios, governed by KL divergence, with validity independent of the stopping rule.
5.5 Sanov’s theorem and likelihood-ratio E-values
Sanov’s theorem describes the exponential rarity scale of empirical distributions under the null; likelihood-ratio evidence inherits this exponential structure via the KL rate. Consequently, LR boundaries scale like $\sqrt{(\log b)/t}$ (moderate deviations), not like $1/b$ (Markov/Ville), completing the connection between the large-deviation geometry of this section and the boundary-selection problem of Remarks˜2.7 and 3.8.
The connection between Sanov’s theorem and E-values can be made precise through the method of types. Fix a distribution $Q$ on $\mathcal{X}$ and consider the likelihood ratio
$E_{n}(Q):=\prod_{i=1}^{n}\frac{Q(X_{i})}{P_{0}(X_{i})}.$
Under $P_{0}$, $\operatorname{\mathbb{E}}_{P_{0}}[E_{n}(Q)]=1$, so $E_{n}(Q)$ is an E-variable. On any sequence with empirical distribution $\hat{P}_{n}=Q$, the log likelihood ratio satisfies
$\frac{1}{n}\log E_{n}(Q)=\sum_{x}Q(x)\log\frac{Q(x)}{P_{0}(x)}=D_{\operatorname{KL}}(Q\|P_{0}).$
Simultaneously, the method-of-types bound implies that under $P_{0}$, the probability of observing type $Q$ is of order $\exp(-nD_{\operatorname{KL}}(Q\|P_{0}))$ up to polynomial factors in $n$. Sanov’s theorem extends this type-based bound from single types to sets of distributions by optimizing $D_{\operatorname{KL}}(Q\|P_{0})$ over the set. The same KL divergence that makes the observation rare under $P_{0}$ (Sanov) also makes the evidence against $P_{0}$ grow rapidly (E-value).
5.6 Duality: Sanov versus inverse Sanov (Bayesian LDP)
Theorem 5.7 (Inverse Sanov).
Level: (L1)++(L2); bridge: Bayes’ rule (posterior given data).
Let $\mathcal{X}$ be a finite set. Let $\pi$ be a prior distribution on $\mathcal{P}(\mathcal{X})$ with $\mathrm{supp}(\pi)\subseteq\mathcal{P}(\mathcal{X})$, and let $\pi_{n}$ denote the posterior given i.i.d. observations $X_{1},\ldots,X_{n}$. Condition on the event $\{\hat{P}_{n}\to\mu\}$ for a fixed $\mu\in\mathrm{supp}(\pi)$. Then $\pi_{n}$ satisfies a large deviation principle on $\mathcal{P}(\mathcal{X})$ restricted to $\mathrm{supp}(\pi)$ with rate function $D_{\operatorname{KL}}(\mu\|\cdot)$ [16]. Precisely, for Borel sets $A\subseteq\mathcal{P}(\mathcal{X})$,
$\limsup_{n\to\infty}\frac{1}{n}\log\pi_{n}(\bar{A})$ $\leq-\inf_{\nu\in\bar{A}\cap\mathrm{supp}(\pi)}D_{\operatorname{KL}}(\mu\|\nu),$ (5.4)
$\liminf_{n\to\infty}\frac{1}{n}\log\pi_{n}(A^{\circ})$ $\geq-\inf_{\nu\in A^{\circ}\cap\mathrm{supp}(\pi)}D_{\operatorname{KL}}(\mu\|\nu).$ (5.5)
The rate function is $\infty$ outside $\mathrm{supp}(\pi)$.
Remark 5.8 (Duality principle).
[…]
Further examples and extensions—including sequential large-deviation arguments, the biased-die illustration, moment constraints/entropic tilting, and simulation studies—are collected in Appendix˜C.
5.7 Process-level implication
Because Bayesian marginal likelihoods are mixtures of likelihood-ratio processes (Proposition˜4.9), the resulting marginal evidence process lies in the process evidence class $\mathcal{E}^{\mathrm{proc}}(H_{0})$ (Definition˜4.1). Hence information-regularized inference inherits anytime validity (Theorem˜4.5) as a structural consequence of its supermartingale representation. The Gibbs posterior is the unique composition law respecting both Proposition˜2.3 and Proposition˜4.3.
6 Exchangeability as symmetry; conformal e-prediction as invariance testing
Exchangeability is a symmetry property at (L2): invariance under the symmetric group action on (L1). Testing it via conformal e-prediction constructs E-values at (L3) valid under the symmetry hypothesis.
6.1 Permutation action and exchangeability
Definition 6.1 (Exchangeability as group invariance).
Let $S_{n}$ denote the symmetric group on $\{1,\ldots,n\}$, acting on $\mathcal{X}^{n}$ by permuting coordinates. For $\pi\in S_{n}$, define
$\pi\cdot(x_{1},\ldots,x_{n}):=(x_{\pi(1)},\ldots,x_{\pi(n)}).$
A law $P$ on $\mathcal{X}^{n}$ is exchangeable if it is invariant under this action:
$(X_{1},\dots,X_{n})\overset{d}{=}(X_{\pi(1)},\dots,X_{\pi(n)})\quad\text{for all }\pi\in S_{n}.$
An infinite sequence $X_{1},X_{2},\ldots$ is exchangeable if every finite marginal $(X_{1},\ldots,X_{n})$ is exchangeable.
Finite exchangeability does not imply an i.i.d. mixture representation. De Finetti’s theorem provides this connection for infinite sequences.
Theorem 6.2 (de Finetti’s representation theorem).
Level: (L2); bridge: permutation invariance $\to$ mixture of i.i.d. laws.
An infinite sequence $X_{1},X_{2},\ldots$ taking values in a standard Borel space $\mathcal{X}$ is exchangeable if and only if there exists a probability measure $\mu$ on the space of probability measures on $\mathcal{X}$ such that, conditional on drawing $P$ from $\mu$, the sequence $(X_{i})$ is i.i.d. from $P$:
$\operatorname{\mathbb{P}}(X_{1}\in A_{1},\ldots,X_{n}\in A_{n})=\int\prod_{i=1}^{n}P(A_{i})\,\mu(dP)$
for all $n$ and all measurable sets $A_{1},\ldots,A_{n}$.
The mixing measure $\mu$ is the directing random distribution identified with the tail $\sigma$-field.
6.2 Conformal e-prediction: testing invariance with E-values
A conformal E-predictor is a function that, given a training sequence $z_{1},\ldots,z_{n}$ and a test object $x$, produces an E-value for each potential label $y$. The E-value measures how “non-conforming” the label $y$ would be for object $x$ given the training data.
Formally, a conformal E-predictor is based on a nonconformity e-measure: a function that maps any sequence $(z_{1},\ldots,z_{m})$ of labeled examples to scores $(\alpha_{1},\ldots,\alpha_{m})$ with average at most 1, equivariant under permutations [41].
Proposition 6.3 (Validity under exchangeability).
Level: (L3); bridge: permutation equivariance of nonconformity scores.
For any conformal E-predictor $f$ and any $n$, if $Z_{1},\ldots,Z_{n},(X,Y)$ are exchangeable, then
$\operatorname{\mathbb{E}}\big[f(Z_{1},\ldots,Z_{n},X,Y)\big]\leq 1.$
That is, the conformal E-predictor produces a valid E-value for the true label under the exchangeability hypothesis.
The proof relies on the permutation equivariance of the nonconformity e-measure: under exchangeability, each of the $(n+1)!$ orderings of the augmented sample $(Z_{1},\ldots,Z_{n},(X,Y))$ is equally likely, and the averaging property ensures that the expected score for the true label is at most 1.
Remark 6.4 (Interpretation as invariance testing).
Conformal e-prediction provides an invariance test: it constructs an E-value that is valid under the symmetry (exchangeability) hypothesis and inflates when symmetry is violated. Large E-values indicate non-conformity, evidence that the new observation $(X,Y)$ does not conform to the pattern established by $Z_{1},\ldots,Z_{n}$ under exchangeability.
6.3 Conformal prediction versus conformal e-prediction
Conformal e-prediction replaces $p$-values with E-values: cross-conformal E-predictors remain valid without the factor-of-two correction, and sequential E-values combine multiplicatively for online exchangeability testing (Proposition˜2.3).
7 Kolmogorov complexity, universal prediction, and gambling
7.1 Prefix complexity and the coding–probability boundary
Definition 7.1 (Prefix Kolmogorov complexity).
Let $U$ be a fixed universal prefix-free Turing machine. The (prefix) Kolmogorov complexity of a finite binary string $x$ is
$K(x)\;:=\;\min\{\,|p|:U(p)=x\,\}.$
The value of $K(x)$ depends on the choice of $U$, but only up to an additive constant (invariance theorem). Consequently, statements of the form $K(x)\geq|x|-c$ are machine-robust.
A central bridge between algorithmic complexity and probability is given by the universal a priori semimeasure $m$ (Solomonoff–Levin), which satisfies
$K(x)\;=\;-\log_{2}m(x)\;+\;O(1),$ (7.1)
so that $K(\cdot)$ coincides with a log-probability only through this universal construction (up to constants). Equation (7.1) is the canonical (and essentially unique) mechanism by which a coding object at level (L4) descends to a probabilistic object at level (L2) [28].
Remark 7.2 (Boundary discipline).
One must not conflate $K(x)$ with $\ell_{P}(x)$ without stating the semimeasure/universality conditions and $O(1)$ constants. In particular, $K(x)\neq-\log P(x)$ for any fixed statistical model $P$; the identity holds only for $m$, never through an arbitrary model. These connections terminate at the computability boundary formalized in §8.3.
7.2 Complexity, compression, and gambling
The connection between prediction, compression, and betting is classical: predictive structure enables improved compression, and improved compression corresponds to systematic growth of gambling capital [8, 9].
In sequential prediction, a predictor $q_{t}(\cdot\mid x^{t-1})$ induces a cumulative log-loss $\sum_{t=1}^{n}-\log q_{t}(X_{t}\mid X^{t-1})$; equivalently, it induces a capital process via the likelihood ratio form. This viewpoint was formalized in Cover’s work on universal portfolios and universal prediction: a universal strategy achieves performance close to the best strategy in hindsight, with regret growing only logarithmically (or sublinearly) in time. Kelly’s result gives the gain in log-growth from side information as $I(X;Y)$ [26].
One algorithmic lens on this phenomenon is through conditional (function) complexity. For strings $x,y$ define
$K_{f}(y\mid x)\;:=\;\min\{\,|p|:U(p,x)=y\,\},$
the length of the shortest program transforming $x$ into $y$. This captures the intrinsic descriptive complexity of the input–output relation and makes precise the slogan “good prediction $\leftrightarrow$ compressibility $\leftrightarrow$ capital growth.”
The normalized maximum likelihood (NML) distribution [38],
$P_{\text{NML}}(x)=\frac{\max_{\theta}L(\theta;x)}{\sum_{x^{\prime}}\max_{\theta}L(\theta;x^{\prime})},$
achieves minimax optimal regret for prediction over the model class. The NML code length admits the decomposition
$\log P_{\text{NML}}(x)=-\max_{\theta}\log L(\theta;x)+\log\sum_{x^{\prime}}\max_{\theta}L(\theta;x^{\prime}),$
a maximum-likelihood description length plus a model-complexity term, providing the standard route to the Minimum Description Length (MDL) principle [35, 20].
7.3 Model multiplicity and a Jeffreys–Lindley analogue in universal betting
Bayes factors quantify evidence as ratios of marginal likelihoods and therefore incorporate an automatic multiplicity/complexity penalty under diffuse alternatives (the familiar “Occam factor”), which is one mechanism behind Jeffreys–Lindley-type behavior at large $n$ [24, 29].
An analogous effect appears in universal prediction / universal betting when one mixes over model classes. Consider mixing over a countable family indexed by $r\in\{0,1,\dots\}$ (e.g., degrees, VC indices, or structural complexity levels) with prior weights $\pi(r)\propto 1/(r+1)^{2}$ or, for exposition, $\pi(r)\propto 1/(r+1)$ truncated at $n$. Then the resulting capital (E-value) process inherits an additive regret term of order $\log(n)$ reflecting uncertainty about the appropriate complexity level. Concretely, in standard universal-prediction analyses one obtains bounds of the form
$\log S_{n}\;\gtrsim\;\sup_{r}\,\log S_{n}^{(r)}\;-\;c\,\log(n+1),$ (7.2)
where $S_{n}^{(r)}$ is the capital process associated with the $r$th class/strategy and $c>0$ is an absolute constant depending on the chosen mixing weights. The term $c\log(n+1)$ is the universal analogue of a multiplicity penalty: it is the cost of not knowing the appropriate class in advance, and it mirrors the marginal-likelihood penalty that appears in Bayes factors under diffuse alternatives.
Cover’s counting arguments for classification illustrate the same geometry: for linear threshold functions in $\mathbb{R}^{d}$ with $n$ points in general position, the number of realizable labelings satisfies $|F|=2\sum_{k=0}^{d}\binom{n-1}{k}$ [7].
Equation (7.2) highlights that coding/complexity penalties arise naturally inside capital (E-value) processes, but only when the underlying construction is sequentially normalized. This sets up the computational boundary result of §7.4: regret-optimal codes (e.g., NML) need not induce supermartingale-valid evidence without an explicit sequential normalization mechanism.
7.4 The computational boundary: codes do not generally yield E-processes
The coding–probability boundary reflects a structural obstruction, not only a terminological distinction.
Proposition 7.3 (Computational boundary).
[…]
Let $\mathcal{C}$ be a coding scheme assigning code lengths $\ell_{\mathcal{C}}(x^{n})$ to finite strings (in bits), and let $P_{0}$ be a null hypothesis. Define the induced process $E_{n}^{\mathcal{C}}:=P_{0}(X^{n})^{-1}\cdot 2^{-\ell_{\mathcal{C}}(X^{n})}$ (using $\log_{2}$ for code lengths). Then:
  1. (i)
    If $\mathcal{C}$ is a prefix-free code corresponding to a probability measure $Q$ (i.e., $\ell_{\mathcal{C}}(x)=-\log_{2}Q(x)$), then $E_{n}^{\mathcal{C}}=Q(X^{n})/P_{0}(X^{n})$ is a likelihood ratio and hence a valid E-process under $P_{0}$ (Proposition˜3.6).
  2. (ii)
    If $\mathcal{C}$ is the NML code for a model class $\mathcal{M}$, then $E_{n}^{\mathcal{C}}$ is not in general a supermartingale under $P_{0}$: NML minimizes worst-case individual-sequence regret, not expected loss, and its sequential factorization does not yield conditional distributions summing to one at each step.
  3. (iii)
    If $\mathcal{C}$ induces the universal semimeasure $m$, then $m(X^{n})/P_{0}(X^{n})$ is a supermartingale under $P_{0}$ for any computable $P_{0}$, yielding a valid but non-computable E-process.
Proof.
(i) follows directly from Proposition˜3.6: any likelihood ratio
$\prod_{t=1}^{n}\frac{q(X_{t}\mid X^{t-1})}{p_{0}(X_{t}\mid X^{t-1})}$
with $q$ a probability measure is a nonnegative supermartingale under $P_{0}$.
(ii) The NML distribution satisfies $\sum_{x^{n}}P_{\text{NML}}(x^{n})=1$ globally, but its sequential factorization fails. Define the “conditional NML” at step $t$ by
$q_{t}^{\text{NML}}(x_{t}\mid x^{t-1}):=\frac{P_{\text{NML}}(x^{t-1},x_{t},\tilde{x}_{t+1},\ldots,\tilde{x}_{n})}{P_{\text{NML}}(x^{t-1},\tilde{x}_{t},\ldots,\tilde{x}_{n})}$
for any completion $(\tilde{x}_{t+1},\ldots,\tilde{x}_{n})$. This depends on the choice of completion and on the full sample size $n$ through the normalizing constant $\sum_{x^{\prime}}\max_{\theta}L(\theta;x^{\prime})$, which involves a supremum over all possible future data. Hence $q_{t}^{\text{NML}}$ is not $\mathcal{F}_{t-1}$-measurable, $\sum_{x_{t}}q_{t}^{\text{NML}}(x_{t}\mid x^{t-1})\neq 1$ in general, and the induced ratio need not form a supermartingale. The failure is structural: regret optimality (minimax over individual sequences) is a different objective from supermartingale structure (conditional expectation bound at each step).
(iii) By the universality of $m$: for any computable measure $P_{0}$, there exists $c>0$ such that $m(x^{n})\geq c\cdot P_{0}(x^{n})$ for all $x^{n}$. The process $M_{n}:=m(X^{n})/P_{0}(X^{n})$ is a nonnegative supermartingale under $P_{0}$ because $m$ dominates all computable measures up to a constant and can be expressed as a mixture of computable predictors, each contributing a valid likelihood ratio, so the mixture is also a supermartingale. The non-computability follows because computing $m(x)$ requires solving the halting problem. ∎
The proposition establishes a structural obstruction: there is no canonical lifting of arbitrary code lengths into E-processes preserving supermartingale structure, except through the universal semimeasure bridge. Regret optimality does not imply supermartingale validity; in particular, regret optimality under static normalization does not imply supermartingale validity under the natural filtration (see Theorem˜B.10 for a formal impossibility result). The obstruction is filtration-level: regret-optimality does not imply stepwise conditional normalization. The distinction between supermartingale-valid evidence processes and static coding criteria is central in contemporary discussions of E-values and testing [22].
7.5 Dawid’s prequential principle and MDL
Dawid’s prequential principle [11] evaluates a forecaster by its sequential predictive performance: each $q(x_{t}\mid x^{t-1})$ is a genuine conditional distribution, so the ratio against a null yields an E-process at (L3). This sequential calibration perspective connects to Goldstein’s [17] prevision-based coherence: both frameworks derive their validity from conditional structure rather than global optimization.
Table 1: Prequential evaluation vs. MDL: convergence and divergence.
Prequential MDL
Objective Sequential calibration Shortest total description
Code Prequential ML: $\prod_{t=1}^{n}p(x_{t}\mid\hat{\theta}_{t-1})$ NML or two-part code
Well-specified Converges to true model Selects true model
Misspecified Best calibrated forecaster Shortest code in class
E-process? Yes (sequential conditionals) Not in general (Proposition˜7.3)
Proposition 7.4 (Prequential codes yield E-processes).
Let $q_{t}(\cdot\mid x^{t-1})$ be a predictive kernel for each $t\geq 1$ (i.e., $q_{t}(\cdot\mid x^{t-1})$ is a probability distribution on $\mathcal{X}$ for each $x^{t-1}$, and $q_{t}$ is $\mathcal{F}_{t-1}$-measurable). Then the process
$E_{n}:=\prod_{t=1}^{n}\frac{q_{t}(X_{t}\mid X^{t-1})}{p_{0}(X_{t}\mid X^{t-1})}$
is a nonnegative martingale under $P_{0}$, hence an E-process. In particular, the prequential ML code $q_{t}(x_{t}\mid x^{t-1}):=p(x_{t}\mid\hat{\theta}_{t-1})$, where $\hat{\theta}_{t-1}$ is the MLE based on $x^{t-1}$, yields a valid E-process [12].
Proof.
Each ratio $e_{t}:=q_{t}(X_{t}\mid X^{t-1})/p_{0}(X_{t}\mid X^{t-1})$ satisfies $\operatorname{\mathbb{E}}_{P_{0}}[e_{t}\mid\mathcal{F}_{t-1}]=\sum_{x_{t}}q_{t}(x_{t}\mid X^{t-1})=1$, so $(E_{n})$ is a martingale by the same argument as Proposition˜3.6. ∎
The NML code, by contrast, does not factorize into valid predictive kernels, precisely the obstruction of Proposition˜7.3(ii).
[…]
7.6 Complexity, Gibbs regularization, and machine learning
Section 2 established that $\ell_{P}(x^{n})=-\log P(x^{n})$ is code length, log-loss, and evidence functional. KL divergence measures excess codelength: $\operatorname{\mathbb{E}}_{Q}[-\log P(X)]=H(Q)+D_{\operatorname{KL}}(Q\|P)$.
Gibbs posteriors (Proposition˜5.1) are entropy-regularized code-length minimizers whose update coincides with exponential-weights aggregation and PAC-Bayes constructions. While not equivalent theories, they share the same KL geometry. MDL is the zero-temperature limit of Gibbs regularization.
Statistical evidence, entropy-regularized inference, PAC-Bayes bounds, weights, and MDL are not separate paradigms but realizations of the same KL geometry. The complexity boundary reflects the distinction between universal coding and model-relative frameworks.
7.7 Arnold’s stochasticity parameter
V.I. Arnold developed a geometric approach to measuring randomness that complements algorithmic and probabilistic definitions [2]. For an ordered sequence $0<x_{1}<x_{2}<\cdots<x_{T}<L$ on a segment of length $L$, Arnold defined the stochasticity parameter as $R=\sum_{i=1}^{T}d_{i}^{2}$, where $d_{i}$ are consecutive spacings. The bi-normalized parameter $s=T\sum_{i=1}^{T}(d_{i}/L)^{2}$ ranges from $s_{\min}=1$ (equally spaced) to $s_{\max}=T$ (complete clustering), with expected value $\operatorname{\mathbb{E}}[s]=2T/(T+1)$ for random placement.
The distribution of spacings under random placement follows a $\text{Dir}(1,\ldots,1)$ distribution, enabling exact moment calculations. Values of $s$ below $\operatorname{\mathbb{E}}[s]$ indicate repulsion (regularity); values above indicate clustering (attraction). This provides a simple geometric statistic connecting number-theoretic structure to spacing-based randomness diagnostics.
The Arnold cat map [3] illustrates the duality between deterministic chaos and apparent randomness: the map $\Gamma:(x,y)\mapsto(x+y,x+2y)\pmod{1}$ generates trajectories that appear random despite being completely deterministic. Such sequences can pass standard statistical test suites [4], illustrating that apparent randomness (high statistical complexity) differs from intrinsic randomness (high Kolmogorov complexity). A truly Kolmogorov-random sequence has no short description, while a chaotic sequence has a short description (the map and initial conditions) but exhibits statistical properties indistinguishable from random.
7.8 Structural synthesis: codes, evidence, and the typed calculus
[…]
8 Synthesis: a typed evidence calculus
8.1 Master diagram
The threads developed in the preceding sections converge on a single observation: negative log predictive probabilities serve as a universal evidence currency, with different theoretical traditions corresponding to different aspects of this currency.
[…]
Figure 1: Typed evidence calculus. Solid arrows denote exact structural bridges (enumerated in Table˜2); dashed arrows denote connections that hold only under boundary conditions (universality, computability, or non-identifiability constraints; see Section˜8.3). KL divergence governs both rarity (Sanov) and evidence growth, with reversed arguments reflecting directional duality. Level (L3) denotes a validity certificate (Ville/optional stopping), distinct from the representational LR/BF geometry at (L2).
The monoidal bridge (Proposition˜2.3) translates multiplicative probability to additive evidence. Loss differences exponentiate to likelihood ratios and Bayes factors (Propositions˜3.6 and 3.9). The evidence class (Proposition˜4.3) formalizes operations preserving validity. KL divergence governs rarity (Sanov, Theorem˜5.3) and growth (Proposition˜5.4) with reversed-argument duality (Theorem˜5.7). Exchangeability provides symmetry (Definitions˜6.1 and 6.3).
The computational boundary (Proposition˜7.3) completes the picture by delineating which coding objects produce valid E-processes and which do not. Of the four coding constructions (marginal likelihood, NML/MDL, prequential ML, Kolmogorov complexity), only marginal likelihoods and Dawid’s prequential ML code (Section˜7.5) yield E-processes automatically; the NML code fails the sequential factorization requirement, and Kolmogorov complexity connects to (L2) only through the non-computable universal semimeasure $m$. The prequential principle provides the critical bridge: because a prequential forecaster specifies genuine conditional distributions $q(x_{t}\mid x^{t-1})$ at each step, its likelihood ratio against a null is a supermartingale by construction, and the resulting E-process inherits all the compositional guarantees of the evidence class. This distinction between prequential and NML codes (Table˜1) mirrors the broader typed discipline: sequential validity is a property of level (L3), and only coding schemes that factorize into bona fide predictive kernels at (L2) can cross the bridge.
8.2 Bridges and non-bridges
Figure 1 summarizes the typed evidence calculus. Table˜2 lists the structural bridges; the non-bridges follow.
Table 2: Structural bridges of the typed evidence calculus (solid arrows in Figure˜1).
Bridge Level Statement Where proved
[…]
Ville (optional stopping) (L3) $\operatorname{\mathbb{P}}(\sup E_{t}\geq c)\leq 1/c$ Theorem˜3.5
Sanov rate (L1)++(L2) Rarity rate $D_{\operatorname{KL}}(\cdot\|P_{0})$ Theorem˜5.3
[…]
Conformal e-prediction (L3) $\operatorname{\mathbb{E}}[f(Z^{n},X,Y)]\leq 1$ under exchangeability Proposition˜6.3
[…]
8.3 Non-bridges: four demarcations of scope
The typed evidence calculus is useful precisely because it makes its limits explicit. Four non-identifications must be maintained.
1. Computability boundary (L4 $\not\to$ L2).
The equality $K(x)=-\log_{2}m(x)+O(1)$ holds only for the universal semimeasure $m$, up to an additive constant depending on the machine, and under computability assumptions. It does not extend to arbitrary probabilistic models $P$; in particular, $K(x)\neq-\log P(x)$ for any fixed statistical model.
2. Sequential factorization obstruction (L4 $\not\to$ L3).
Not every code length induces a valid E-process. The NML code assigns a well-defined distribution over complete sequences, but its sequential factorization does not yield conditional distributions summing to one at each step (Proposition˜7.3(ii)): the normalizing constant depends on the full sample size and on future data through the supremum over parameters. Prequential codes avoid this obstruction by construction, maintaining genuine predictive kernels at each step (Section˜7.5). This is the typed explanation of why MDL model selection and sequential evidence accumulation, though both grounded in coding theory, require different structural assumptions.
3. Finite vs. infinite exchangeability (L2).
De Finetti’s theorem (Theorem˜6.2) requires an infinite sequence. For finite samples, exchangeability does not imply an i.i.d. mixture representation, and the distinction between exchangeability and the i.i.d. property is genuine, a fact that conformal methods must respect.
4. Misspecification (L2/L3).
The evidence growth rate $D_{\operatorname{KL}}(P_{1}\|P_{0})$ of Proposition˜5.4 assumes the true distribution is $P_{1}$. Under misspecification, when neither $P_{0}$ nor $P_{1}$ is correct, the growth rate becomes $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})$, and the E-process may not accumulate evidence in the expected direction. Developing constructions robust to misspecification while retaining anytime validity is an active area [21].
The preceding sections show that probability, evidence, deviations, entropy-regularized methods, and coding complexity form layers of a single calculus. The four non-bridges above are not defects but structural features: they delineate precisely where probabilistic, sequential, and computational reasoning diverge, and where specific coding theorems (Kraft’s inequality, the prequential principle, the universality of $m$) are needed to cross the boundary.
9 Discussion
Representation versus construction.
The central message of this paper is that likelihood ratios are the canonical representation of sequential evidence within the coherent predictive/log-loss subclass. The Fubini/Bayes-risk decomposition (Theorem˜2.6) identifies the likelihood ratio as the optimal evidence statistic; Markov/Ville inequalities then provide anytime-valid certificates, but they do not by themselves identify Bayes-optimal rejection regions. Cutoffs are execution-theoretic: they depend on losses, priors, and the monitoring protocol (Section˜3.7), not on the evidence representation. In this sense, within the log-loss/Bayes-risk coherent subclass, sequential evidence admits a likelihood-ratio representation; outside this subclass, e-processes may exist without such a representation [see 27, for the general composite-null theory]. This separation—representation at the evidence layer, cutoffs at the decision layer—resolves apparent tensions between Bayesian and frequentist uses of E-values.
Relation to the e-process literature.
General e-processes, defined by optional-stopping control, strictly generalize test (super)martingales and product-of-e-value constructions [33, Def. 7.3]. Our paper does not claim that all e-processes are likelihood ratios or mixtures thereof; rather, we characterize the coherent predictive subclass in which the likelihood-ratio representation is canonical. The typed bridge and non-bridge theorems identify precisely where coding objects, complexity measures, and general validity-only constructions fit—or provably fail to fit—within this coherent framework. We view the broader e-process programme [1, 33] and our typed representation calculus as complementary: the former emphasizes universal validity constructions, the latter emphasizes the structural semantics that distinguish Bayes-risk-optimal evidence from validity certificates alone.
Coding, MDL, and the computational boundary.
The computational boundary theorem (Proposition˜7.3) is a typed non-bridge result: coding objects descend to probability only via specific semimeasure conditions. NML codes fail filtration-measurability; prequential codes succeed because they are genuine predictive distributions; the universal semimeasure succeeds but is non-computable. This is exactly the bridge-versus-non-bridge discipline that the Dawid/Grünwald reconciliation of MDL and Bayesian inference requires.
Open directions.
Several directions remain open: scalable E-process constructions beyond low-dimensional settings; the interplay between computational constraints and algorithmic complexity; robust E-processes under misspecification [21]; exchangeability diagnostics in streaming and high-dimensional regimes; and extending the typed framework to confidence sequences [23, 44] and to sequential monitoring under distribution shift. Our aim is not to replace general e-process constructions, but to clarify the structural conditions under which likelihood ratios provide a coherent evidential representation.
Appendix A Why four levels? An intuitive guide
The typing discipline of Section˜1.2 may appear formal. This appendix offers everyday analogies to motivate why keeping the four levels separate prevents confusion.
(L1) Sequences: the raw recording.
Think of a sequence $x^{n}$ as a recording of $n$ coin flips written on a slip of paper. The slip contains data and nothing else, no model, no interpretation, no claim about fairness. It is a concrete object in the world.
(L2) Laws: the claimed mechanism.
A law $P$ is a claim about the mechanism that generated the recording, for instance, “the coin is fair.” Two different people can hold different claims (different $P$’s) about the same slip. The loss functional $\ell_{P}(x^{n})=-\log P(x^{n})$ measures how surprised a person holding claim $P$ would be upon reading the slip. This is a bridge from (L2) to (L1): it requires both a claim and a recording.
(L3) Processes: the sequential game.
Now imagine reading the slip one entry at a time and placing bets as you go. The running wealth of a bettor is an E-process: it lives on a timeline with a filtration and stopping rule. Ville’s inequality says that no matter when you decide to stop peeking, a fair game cannot make you rich with high probability. Claims about optional stopping, anytime validity, and sequential error control only make sense at this level, they require the notion of “time” and “when to stop,” which a static law or a static recording does not provide.
(L4) Codes: the compression program.
Kolmogorov complexity asks: what is the shortest computer program that prints the recording? This is a question about Turing machines and algorithms, not about probability or betting. The bridge to (L2) is the coding theorem $K(x)=-\log_{2}m(x)+O(1)$, but the “O(1)O(1)” hides a constant that depends on the choice of programming language (universal machine). Forgetting this, treating $K(x)$ as if it were $\log P(x)$ for some fixed statistical model, is like confusing the file size of a JPEG with the probability of the photograph.
Why does this matter in practice?
Conflating levels leads to well-known errors: interpreting $p$-values after optional stopping without correction (mixing L2 and L3); claiming that a short description proves a model is “true” (mixing L4 and L2); or treating exchangeability of a finite sample as if de Finetti’s infinite-sequence theorem applied directly (mixing finite and infinite within L2). The typing discipline is simply a checklist that prevents these conflations.
Appendix B Foundations of the typed evidence system
This appendix formalizes the typed separation used throughout the paper. Its purpose is not to introduce new probability theory, but to distinguish three logically distinct layers: (i) representational geometry (Radon–Nikodým / likelihood-ratio structure), (ii) validity certificates (martingale/Ville control), and (iii) boundary/decision design. Conflating these layers leads to apparent disagreements about whether likelihood ratios or e-processes are “primitive.” The typed system makes explicit that these are different roles, each with its own algebraic and analytic properties.
Concretely, we give self-contained proofs for the structural results invoked in the main text: uniqueness of the log-score map, the convex-algebraic properties of the evidence class, the KL growth-rate theorem, and the computational boundary obstruction.
Remark B.1 (Scope: e-processes vs. test supermartingales).
An e-process is defined by optional-stopping control and strictly generalizes test (super)martingales [33, Def. 7.3]. In particular, product constructions from sequential e-values yield e-processes, but do not exhaust the class. Our representation results apply to the coherent predictive subclass in which evidence admits a log-loss/Bayes-risk interpretation; within this subclass, likelihood-ratio martingales (and mixtures) provide the canonical representation (Theorem˜2.6).
B.1 Uniqueness of the log-score transformation
The central bridge in the main text is the log-score map $\mathcal{L}:P\mapsto-\log P$ (Definition˜2.1). We show that this transformation is essentially determined by structural requirements.
Theorem B.2 (Uniqueness of the additive likelihood transform).
Let $\phi:(0,\infty)\to\mathbb{R}$ be continuous and strictly monotone. Suppose that for all $a,b>0$,
$\phi(ab)=\phi(a)+\phi(b).$
Then $\phi(x)=c\log x$ for some constant $c\neq 0$.
Proof.
The functional equation states that $\phi$ is a continuous homomorphism from the multiplicative group $(\mathbb{R}_{>0},\cdot)$ to the additive group $(\mathbb{R},+)$. Define $\psi:\mathbb{R}\to\mathbb{R}$ by $\psi(t):=\phi(e^{t})$. Then $\psi(s+t)=\phi(e^{s+t})=\phi(e^{s}\cdot e^{t})=\phi(e^{s})+\phi(e^{t})=\psi(s)+\psi(t)$, so $\psi$ satisfies Cauchy’s functional equation. Since $\phi$ is continuous, $\psi$ is continuous, and the only continuous solutions to Cauchy’s equation are $\psi(t)=ct$ for some constant $c$. Therefore $\phi(x)=\psi(\log x)=c\log x$. Strict monotonicity forces $c\neq 0$. ∎
Corollary B.3.
Among continuous order-preserving transforms, the negative logarithm $x\mapsto-\log x$ is the unique map (up to positive scaling) that converts multiplicative likelihood ratios into additive evidence increments.
Proof.
By Theorem˜B.2, any such transform has the form $\phi(x)=c\log x$. Order-reversal (smaller probabilities yield larger evidence) requires $c<0$; normalizing to c=1c=-1 gives $\phi(x)=-\log x$. ∎
Thus the log-score map of Definition˜2.1 is not only conventional; it is structurally determined by multiplicativity and additivity.
B.2 The evidence class as a convex algebra
Let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\operatorname{\mathbb{P}}_{H_{0}})$ be a filtered probability space under $H_{0}$. We collect the algebraic closure properties of the E-process class $\mathcal{E}^{\mathrm{proc}}(H_{0})$ stated in Theorem˜1.1 and prove them in a self-contained manner.
Proposition B.4 (Convex closure).
If $(E^{(1)}_{t})$ and $(E^{(2)}_{t})$ are E-processes under $H_{0}$ and $\lambda\in[0,1]$, then
$E_{t}:=\lambda E^{(1)}_{t}+(1-\lambda)E^{(2)}_{t}$
is an E-process under $H_{0}$.
Proof.
Nonnegativity is clear. By linearity of conditional expectation,
$\operatorname{\mathbb{E}}_{H_{0}}[E_{t}\mid\mathcal{F}_{t-1}]=\lambda\,\operatorname{\mathbb{E}}_{H_{0}}[E^{(1)}_{t}\mid\mathcal{F}_{t-1}]+(1-\lambda)\,\operatorname{\mathbb{E}}_{H_{0}}[E^{(2)}_{t}\mid\mathcal{F}_{t-1}]\leq\lambda\,E^{(1)}_{t-1}+(1-\lambda)\,E^{(2)}_{t-1}=E_{t-1}.$
Hence $(E_{t})$ is a nonnegative supermartingale with $\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$. ∎
Proposition B.5 (Predictable stopping).
Let $(E_{t})$ be an E-process under $H_{0}$ and $\tau$ a stopping time with respect to $(\mathcal{F}_{t})$. Then $(E_{t\wedge\tau})_{t\geq 0}$ is an E-process under $H_{0}$.
Proof.
By the optional stopping theorem for nonnegative supermartingales, $(E_{t\wedge\tau})$ is a nonnegative supermartingale. Since $E_{0\wedge\tau}=E_{0}$, the initial condition $\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$ is preserved. ∎
Proposition B.6 (Countable stitching).
Let $\{(E^{(k)}_{t})\}_{k\in\mathbb{N}}$ be E-processes under $H_{0}$ and $\{\lambda_{k}\}_{k\in\mathbb{N}}$ nonnegative weights with $\sum_{k}\lambda_{k}\leq 1$. Then
$E_{t}:=\sum_{k=1}^{\infty}\lambda_{k}E^{(k)}_{t}$
is an E-process under $H_{0}$.
Proof.
Nonnegativity is clear. By the monotone convergence theorem and the supermartingale property of each $(E^{(k)}_{t})$,
$\operatorname{\mathbb{E}}_{H_{0}}[E_{t}\mid\mathcal{F}_{t-1}]=\sum_{k}\lambda_{k}\,\operatorname{\mathbb{E}}_{H_{0}}[E^{(k)}_{t}\mid\mathcal{F}_{t-1}]\leq\sum_{k}\lambda_{k}\,E^{(k)}_{t-1}=E_{t-1}.$
The initial condition follows from $\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]=\sum_{k}\lambda_{k}\,\operatorname{\mathbb{E}}_{H_{0}}[E^{(k)}_{0}]\leq\sum_{k}\lambda_{k}\leq 1$. ∎
These results establish that the class of E-processes forms a convex set, closed under scaling by $c\in(0,1]$, predictable stopping, and countable convex mixtures, as stated in Theorems˜1.1 and 4.3. Note that this class is not a cone in the algebraic sense: multiplication by constants greater than one generally violates the unit-expectation constraint $\operatorname{\mathbb{E}}_{H_{0}}[E_{0}]\leq 1$, so only scaling by $c\in(0,1]$ preserves membership. These convex-algebraic properties are compatible with the fork-convexity geometry studied in the betting and e-process literature [33, Sec. 7.2].
Proposition B.7 (Layer separation).
The following are logically distinct objects:
  1. 1.
    A likelihood-ratio representation $E=dQ/dP$ (representation layer);
  2. 2.
    A supermartingale property $\operatorname{\mathbb{E}}_{P}[E_{t}\mid\mathcal{F}_{t-1}]\leq E_{t-1}$ (validity layer);
  3. 3.
    A boundary rule $\tau=\inf\{t:E_{t}\geq b\}$ (decision layer).
Each may be specified independently of the others, and optimality statements at one layer do not imply optimality at another. In particular, a process may possess a likelihood-ratio representation without being a supermartingale (if the denominator is misspecified), and a valid e-process may admit no likelihood-ratio representation at all.
Proof.
Independence of (1) and (2): any ratio $dQ/dP$ is a $P$-martingale when $Q\ll P$, but ratios formed under a misspecified model (i.e., $Q$ not absolutely continuous on the filtration) need not be supermartingales. Conversely, e-processes constructed via convex mixtures (Propositions˜B.4 and B.6) need not admit a single likelihood-ratio representation. Independence of (3): the boundary $b$ is a free parameter that controls the type-I/power trade-off but does not affect validity or the representational structure of $(E_{t})$. ∎
Remark B.8 (Nonparametric validity).
Uniform validity over a null class $\mathcal{P}$ is a statement about the certificate (validity) layer: $\sup_{P\in\mathcal{P}}\operatorname{\mathbb{E}}_{P}[E_{t}\mid\mathcal{F}_{t-1}]\leq E_{t-1}$. It does not constrain the alternative-side geometry of $\log E_{t}$, which may inherit KL/MDP structure when a likelihood-ratio representation exists (Proposition˜B.9).
B.3 KL divergence as growth-rate functional
We give a self-contained proof of the growth-rate characterization stated in Proposition˜5.4.
Proposition B.9 (KL growth rate — formal statement).
Let $P_{1}$ and $P_{0}$ be probability measures on a measurable space $(\mathcal{X},\mathcal{B})$ with $P_{1}\ll P_{0}$ and $D_{\operatorname{KL}}(P_{1}\|P_{0})<\infty$. Let $X_{1},X_{2},\ldots$ be i.i.d. under $P_{1}$ and define the likelihood-ratio E-process
$E_{n}:=\prod_{i=1}^{n}\frac{dP_{1}}{dP_{0}}(X_{i}).$
Then
$\frac{1}{n}\log E_{n}\;\xrightarrow{\;\mathrm{a.s.}\;}\;D_{\operatorname{KL}}(P_{1}\|P_{0})\quad\text{under }P_{1}.$
Under misspecification, if $X_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P_{\mathrm{true}}$ with $P_{\mathrm{true}}\ll P_{0}$ and $P_{\mathrm{true}}\ll P_{1}$, then
$\frac{1}{n}\log E_{n}\;\xrightarrow{\;\mathrm{a.s.}\;}\;D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1}).$
Proof.
Write
$\tfrac{1}{n}\log E_{n}\;=\;\tfrac{1}{n}\sum_{i=1}^{n}\log\frac{dP_{1}}{dP_{0}}(X_{i}).$
Under $P_{1}$ the summands are i.i.d. with mean $D_{\operatorname{KL}}(P_{1}\|P_{0})$; the strong law gives the first claim. Under $P_{\mathrm{true}}$ the mean is
$\operatorname{\mathbb{E}}_{P_{\mathrm{true}}}\!\Bigl[\log\frac{dP_{1}}{dP_{0}}(X)\Bigr]$ $=\operatorname{\mathbb{E}}_{P_{\mathrm{true}}}\!\Bigl[\log\frac{dP_{\mathrm{true}}}{dP_{0}}(X)\Bigr]-\operatorname{\mathbb{E}}_{P_{\mathrm{true}}}\!\Bigl[\log\frac{dP_{\mathrm{true}}}{dP_{1}}(X)\Bigr]$
$=D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1}),$
and the strong law gives the second claim. ∎
The misspecification formula shows that the E-process drifts downward (evidence against $H_{1}$) precisely when $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})>D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})$, i.e., when the true distribution is closer to $P_{0}$ than to $P_{1}$ in KL divergence.
B.4 Computational boundary: an impossibility result
We formalize the obstruction between coding objects and supermartingale-valid evidence stated in Proposition˜7.3.
Theorem B.10 (No canonical lifting of static codes).
Let $\ell:\mathcal{X}^{n}\to\mathbb{R}_{\geq 0}$ be a code-length function and $P_{0}$ a null hypothesis with predictive kernel $p_{0}(\cdot\mid x^{t-1})$. Define the induced process
$E_{t}:=\frac{\exp(-\ell(X^{t}))}{P_{0}(X^{t})},\qquad t=1,\dots,n.$
Suppose $(E_{t})_{t=0}^{n}$ is a supermartingale under $P_{0}$ with $E_{0}=1$. Then $\ell$ must factorize sequentially: there exist functions $\ell_{t}:\mathcal{X}^{t}\to\mathbb{R}_{\geq 0}$ with $\sum_{x_{t}}p_{0}(x_{t}\mid x^{t-1})\exp(-\ell_{t}(x^{t})+\ell_{t-1}(x^{t-1}))\leq 1$ for all $x^{t-1}$, and $\ell(x^{n})=\sum_{t=1}^{n}[\ell_{t}(x^{t})-\ell_{t-1}(x^{t-1})]$.
Equivalently, $\exp(-\ell)$ must arise from a sequentially normalized predictive distribution $q_{t}(\cdot\mid x^{t-1})$ satisfying $\sum_{x_{t}}q_{t}(x_{t}\mid x^{t-1})=1$, via
$\ell(x^{n})=-\sum_{t=1}^{n}\log q_{t}(x_{t}\mid x^{t-1}).$
In particular, static regret-optimal codes (e.g., NML) whose normalizing constant depends on the full sample size $n$ do not in general factorize in this way and hence do not induce E-processes.
Proof.
The supermartingale condition $\operatorname{\mathbb{E}}_{P_{0}}[E_{t}\mid\mathcal{F}_{t-1}]\leq E_{t-1}$ gives
$\sum_{x_{t}}p_{0}(x_{t}\mid X^{t-1})\,\frac{\exp(-\ell(X^{t-1},x_{t}))}{P_{0}(X^{t-1},x_{t})}\;\leq\;\frac{\exp(-\ell(X^{t-1}))}{P_{0}(X^{t-1})},$
where $\ell(X^{t-1})$ denotes the marginal code length at time t1t-1. Since $P_{0}(X^{t-1},x_{t})=P_{0}(X^{t-1})\,p_{0}(x_{t}\mid X^{t-1})$, the $p_{0}$ terms cancel:
$\sum_{x_{t}}\frac{\exp(-\ell(X^{t-1},x_{t}))}{\exp(-\ell(X^{t-1}))}\;\leq\;1.$
Define $q_{t}(x_{t}\mid X^{t-1}):=\exp(-\ell(X^{t-1},x_{t})+\ell(X^{t-1}))$. The inequality above states $\sum_{x_{t}}q_{t}(x_{t}\mid X^{t-1})\leq 1$, so $q_{t}(\cdot\mid X^{t-1})$ is a sub-probability kernel. If equality holds at each step (the martingale case), $q_{t}$ is a genuine predictive distribution and $\ell(x^{n})=-\sum_{t=1}^{n}\log q_{t}(x_{t}\mid x^{t-1})$.
For NML, the normalizing constant $C_{n}:=\sum_{x^{\prime}\in\mathcal{X}^{n}}\max_{\theta}L(\theta;x^{\prime})$ depends on the full sample size $n$. The “conditional NML” at step $t$,
$q_{t}^{\mathrm{NML}}(x_{t}\mid x^{t-1}):=\frac{P_{\mathrm{NML}}(x^{t-1},x_{t},\tilde{x}_{t+1},\ldots,\tilde{x}_{n})}{P_{\mathrm{NML}}(x^{t-1},\tilde{x}_{t},\ldots,\tilde{x}_{n})},$
depends on the choice of completion $(\tilde{x}_{t+1},\ldots,\tilde{x}_{n})$ and on future data through $C_{n}$. Hence $q_{t}^{\mathrm{NML}}$ is not $\mathcal{F}_{t-1}$-measurable and does not sum to one over $x_{t}$, violating the sequential normalization requirement. ∎
Corollary B.11.
Regret optimality under static normalization does not imply supermartingale validity under the natural filtration.
This establishes a structural boundary between coding-theoretic optimality and sequential evidence validity, as stated in Proposition˜7.3.
Lemma B.12 (Likelihood-ratio martingale and reciprocal).
Let $P_{0},P_{1}$ be two measures with $P_{1}\ll P_{0}$ and define $B_{n}=dP_{1}/dP_{0}(X^{n})$. Then $(B_{n})_{n\geq 0}$ is a nonnegative $P_{0}$-martingale with $\operatorname{\mathbb{E}}_{0}[B_{n}]=1$. If also $P_{0}\ll P_{1}$, then $1/B_{n}=dP_{0}/dP_{1}(X^{n})$ is a nonnegative $P_{1}$-martingale with $\operatorname{\mathbb{E}}_{1}[1/B_{n}]=1$.
Proof.
Both claims follow from the tower property applied to the Radon–Nikodym derivative of one measure with respect to the other, restricted to the filtration generated by $X^{n}$. ∎
For general e-process constructions beyond product-of-stepwise e-values, see Ramdas [33, Ch. 7], including the distinction between test supermartingales and e-processes and several general constructions for composite testing problems.
Bibliographical note
The distinction between test (super)martingales (including product-form constructions) and the more general notion of e-processes is emphasized in Ramdas [33, Def. 7.3], which also records the nontrivial equivalence between stopping-time formulations and domination by test martingale families (proved via Snell envelopes and Doob decompositions in Ramdas et al. 34). See Ramdas [33, Sec. 7.2–7.4 and bibliographical note] for historical context and for examples (e.g. exchangeability testing) where e-processes exhibit behavior not captured by test supermartingales; geometric closure notions such as fork-convexity play a key role. Our convex-algebraic closure properties (Section˜B.2) are compatible with, and complementary to, the fork-convexity geometry emphasized in the betting literature.
Appendix C Rate theory: extended examples and sequential arguments
This appendix collects supplementary material on large-deviation rate theory, concrete examples, and simulation studies that complement the main-text development in Section 5.
C.1 Sequential extensions
The multiplicative structure of E-values enables sequential large-deviations analysis. Define a running E-process
$E_{n}=\prod_{i=1}^{n}e_{i},\qquad\operatorname{\mathbb{E}}_{P_{0}}[e_{i}\mid X_{1},\ldots,X_{i-1}]\leq 1,$
and apply supermartingale arguments: for any stopping time $\tau$,
$\operatorname{\mathbb{P}}_{P_{0}}(E_{\tau}\geq 1/\alpha)\leq\alpha.$
Sanov’s theorem provides the static (fixed-$n$) large-deviations picture. The E-value framework extends this to a dynamic, sequential picture. Recent work has shown that sequential versions of Sanov’s theorem can be derived directly from the theory of E-processes, and that the KL divergence rate function appears as the almost-sure growth rate of the optimal E-process [21].
C.2 A concrete example: the biased die
Consider a biased die with outcome probabilities $Q=(1/3,2/3)$ on $\{0,1\}$. Suppose we observe an empirical mean $\frac{1}{n}\sum_{i=1}^{n}X_{i}=3/4$, corresponding to empirical distribution $\hat{P}=(1/4,3/4)$.
Sanov’s theorem gives $Q^{n}(E_{n})\approx\exp(-n\,D_{\operatorname{KL}}(\hat{P}\|Q))$, where
$D_{\operatorname{KL}}(\hat{P}\|Q)=\frac{1}{4}\log\frac{1/4}{1/3}+\frac{3}{4}\log\frac{3/4}{2/3}=\frac{1}{4}\log\frac{3}{4}+\frac{3}{4}\log\frac{9}{8}.$
From the E-value perspective, the likelihood-ratio E-value testing $H_{0}:P=Q$ with alternative $\hat{P}$ grows exponentially at rate $D_{\operatorname{KL}}(\hat{P}\|Q)$. The same KL divergence that makes the observation rare under $Q$ (Sanov) also makes the evidence against $Q$ grow rapidly (E-value), hence the duality between large deviations and hypothesis testing.
C.3 Moment constraints and entropic tilting
The same variational structure appears when inference is conducted under moment constraints. Let $\mathcal{M}$ denote the set of distributions satisfying $\operatorname{\mathbb{E}}_{P}[g(X)]=m$. Then the entropy-minimizing solution
$P^{\star}=\operatorname{arg\,min}_{P\in\mathcal{M}}D_{\operatorname{KL}}(P\|P_{0})$
has exponential-tilted (Gibbs) form.
This construction can be viewed as an L1–L2 composition mediated by KL regularization: partial empirical information (moment values, L1) is fused with a baseline model ($P_{0}$, L2) to yield coherent predictive objects that generate E-processes under the null. See Polson and Zantedeschi [32] for the moment-constrained inference framework and its large-deviation analysis, which develops a curvature-adaptive version of exchangeable updating under finite constraints.
C.4 Simulation: evidence growth under correct and misspecified alternatives
Figure 2: Simulated paths of $\log E_{n}$ for Bernoulli likelihood-ratio E-processes with null $P_{0}=\mathrm{Bern}(0.5)$. (a) Correct specification: data from $P_{1}=\mathrm{Bern}(0.65)$; E-process grows at rate $D_{\operatorname{KL}}(P_{1}\|P_{0})\approx 0.046$ (dashed line), confirming Proposition˜5.4. (b) Misspecification: data generated from $P_{\mathrm{true}}=\mathrm{Bern}(0.55)$ but the E-process uses $P_{1}=\mathrm{Bern}(0.80)$; the effective growth rate is $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})-D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})\approx-0.154$ (dashed line), and the E-process drifts downward, it accumulates evidence for the null despite the null being false.
Figure 2 illustrates the contrast between correct specification and misspecification of the alternative in a Bernoulli setting. Under correct specification (panel a), the five sample paths of $\log E_{n}$ cluster tightly around the theoretical growth rate $n\cdot D_{\operatorname{KL}}(P_{1}\|P_{0})$, confirming Proposition˜5.4. Under misspecification (panel b), the alternative $P_{1}=\mathrm{Bern}(0.80)$ is too extreme relative to the true $P_{\mathrm{true}}=\mathrm{Bern}(0.55)$: the KL penalty $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{1})$ exceeds the KL gain $D_{\operatorname{KL}}(P_{\mathrm{true}}\|P_{0})$, and the net rate is negative. The E-process shrinks, and the sequential test never rejects, despite the null being false.
Robust constructions mixing over alternatives (Section˜3.5) mitigate this sensitivity.
Remark C.1 (Construction classes versus representation theorems).
Our LR representation results are representation statements (what evidence must look like under coherence / Bayes-risk structure), not claims that all useful e-processes must be built as products of one-step e-values. General e-process constructions—including those not representable as multiplicative test martingales for certain composite or nonparametric problems—are discussed in the game-theoretic safe-testing literature; see, e.g., Ram [1]. The point here is orthogonal: whenever a coherent predictive state exists, the induced evidence process is an LR (or mixture of LRs), and boundary choice is governed by Bayes risk rather than Markov calibration.
Acknowledgments
We thank Aaditya Ramdas and Nick Koning for detailed comments on an earlier version of the manuscript.